A Complete Bibliography of AT&T Technical Journal, 1990–1996, and Bell Labs Technical Journal, 1996–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

16 March 2020 Version 1.11

Title word cross-reference

1990s [Bar94, HJKN95]. 1A [DMM^{+}95]. 1B [BGM^{+}95, FO95].

2 [BBD^{+}97, YS95]. 2000 [DJ92, GH97, HW98, HHJ^{+}94, HBCS93, MASY91, TSS95]. 2001 [CB95, Rab95]. 21st [Mid97]. 2500 [EKD93].

4ESS [ACFM94]. 4ESS^{T M} [GHJ98].

5ESS [BFF^{+}91, Far92, GHK^{+}93, HHJ^{+}94, HBCS93, LMN^{+}92, TSS95, WMS95]. 5ESS-2000 [HHJ^{+}94, HBCS93, TSS95]. 5ESS(R) [BF97, GH97, KBP97]. 5ESS(R)-2000 [GH97].

7 [WJC99].

9-1-1 [MJP^{+}96]. 95 [CWH^{+}96].

AAL [BBD^{+}97]. AAL-2 [BBD^{+}97]. accelerated [LCK^{+}98]. Access [RMR^{+}97, WJC99, ALZ96, ALS99, BBR^{+}96, BBH^{+}99, CCJ^{+}99, DDM^{+}96, DDJ^{+}99, FS96, FSW98, GP96a, GMSW96, KRZ98, LSWZ96, SSM94, SK96, Ufo99, ZTL96]. accounting [BMB^{+}99]. achieve [Far93]. Achieving [BCG^{+}94]. acoustic [BMHS98, EET95]. across [LMMT95, SR98]. Active [BHZ97]. adaptation [BCEN99, BBD^{+}97]. adaptive [HMH99, MK98, RM97]. add [GS99]. add/drop [GS99]. added [ZN99]. ADEPT [Sch98]. admission [HW98]. Advanced [JN99, SSW93, SR98, Sch98]. Advances [EP97, JPT95]. Adventures [Tow99]. Advice [Tot91]. against [CLW95]. age [Ros97]. agent [BDGH95]. agents [BBB^{+}99]. Aided [SKS91]. Algorithmic [DS90, ULS90]. Algorithms [BH91, Hol90, ABD^{+}98, MZ97]. allocation [ABD^{+}98]. Alternative [HZ97]. altitudes [LTM97]. AM [McG92]. American [GS97]. amplified [LMW95, MJSS95]. Amplifiers [ZGSD92, SSZS99]. Analysis [IPU91, AKM94, Dem99, DDJ^{+}99, GHJ98, HJ97, HS98, LB97, TB98, Ufo99, Yu98]. anomaly [HMH99]. antennas [BKcL^{+}99, Fos96]. AnyMedia(R) [CCJ^{+}99]. Application [FM95, SD90, Tot91, BKR^{+}93, ZMRR95, vSV97]. Applications [BH91, FKF90, HRRS90, RRW91, VPR^{+}90, WKS^{+}91, AAE^{+}93, BCC^{+}95, DK93a, EA97, FE91, FD90, JC95, MK98, RMN97, SSM94]. applied [GH97]. Applying [CLG92]. Approach [HMD^{+}95, MC91, AKM94, FJ94, HMH99, Nag98, Spr98, YLB96, Yu98]. approximation [Kha97]. Aqueous [WDM^{+}92]. Architectural [FS96]. Architecture [BCF^{+}96, FM95, LB97, WBE^{+}91, ASSS96, BRSS97, BBH^{+}99, Fos96, GHN^{+}99, KNS^{+}92, KK98, LPS98, PR99, RHNS91, WJC99, KDF90]. Architectures [BDR^{+}92, ALZ96, AMRMV99, Che98, CS98, DDHQ98, KJPR97, PLS90, Ufo99]. area [DHVS^{+}98, RZ99]. art [HHMP94]. art/craft [HHMP94]. ASCC [DBL96]. ASN.1 [Mit94]. Assembly [AHM^{+}92, ABS92, BJL92, Ell92, LRW90]. asset [BDNT96]. assets [LTM97]. Assistant [HMMS90]. assuring [FJ94]. Asynchronous [CCS95, Sub91]. AT&T [AL95, BEF^{+}95, BCD^{+}94, BLR95, BD95, BPP92, BFF^{+}91, CK94, ELC92, FE91, FE92, FSV91, Fos91, GP96b, HCWW91, HS95, KNS^{+}92, PLS90, Sei95, SSM94, VPR^{+}90, WMS95, WKS^{+}91, YFF94]. ATM [ANN97, BBD^{+}97, BCC^{+}95, BF97, CKLM98, CT97a, CT97b, Che98, CDT98, DDM^{+}96, EP97, HW98, KRZ98, Noh98, SH99, TWN95, Vee97, YFF94]. ATM-based [BCC^{+}95]. ATT [HK95]. ATT-GIS [HK95]. attributes [WMS95]. Audio [BEF^{+}95, BF90, JC95, JLP90, Rab95]. audits [FHM^{+}91]. Augsburg [HK95]. Automated [ABS92, Sub91, Hol90]. Automatic [WGHL93, JPT95].

BALI [EKK^{+}97]. band

[KM97, KRZ98, Tuc93]. bands [BDK96]. bandwidth [ABD^{+}98, GHN^{+}99, ABDH99]. banking [JSS^{+}95]. Barker [BDK96]. base [LB97]. Based [BF90, LGPR93, Mil93, SD90, BDGH95, BMW94, BCC^{+}95, FCT^{+}99, HLMS93, HSGR99, LTM97, LB97, RM97, SW99]. basic [BCRV95]. beat [GHJ98]. Bell [Ano96a, Ano96b, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Spr98]. Benefits [KSS93]. Benign [MVC90]. Best [DMM^{+}95, CC96, GUP98, YLB96, vSV97]. best-in-class [GUP98]. Better [DMM^{+}95, MG91]. beyond [Rex91]. billing [BMB^{+}99]. bitstream [ZYB98]. block [ABDH99]. Board [AHM^{+}92]. Boundary [YBC^{+}94]. Boundary-scan [YBC^{+}94]. BRI [ABI^{+}93a]. bridge [HLMS93]. Bridging [SS98]. briefs [Ano96a, Ano96b, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d]. bring [TWN95]. Bringing [GP96b]. Broadband [BBH^{+}99, DK93a, LSWZ96, RMR^{+}97, ASSS96, DDM^{+}96, GHN^{+}99, KRZ98, WWS^{+}99, ZTL96]. build [vdBC97]. Building [FHO92, ABDH99, CS98, vSV97]. bullet [DBL96, Gel94]. business [ALZ96, BDPU94, CS96, GP96b, KRS98, Kyt96, SS91, vSV97]. businesses [SW95].

Cable [ABS92, KGC95, SAES95, ZMRR95]. Cabling [HE93]. CAD [Fos91]. CAE [Fos91]. CAE/CAD [Fos91]. Call [HHP91, Che98, HJKN95, PSS96]. cancellation [BMHS98]. Capabilities [BCM91, LMN^{+}92]. Capability [FF92, LMN90]. capable [Bus98]. capacity [Chr99, FKCR92, KD97, ZN99]. Card [SSM94]. carrier [FSW98, KBB99]. carrier-scale [FSW98]. carriers [ZN99]. case [CFH^{+}99, FLJL^{+}98, KK98, CLG92]. CATV [WWS^{+}99]. cause [Yu98]. CDMA [BF97, EKK^{+}97, Pit99, ZN99]. cdma2000 [KLR98]. cell [BLWW94]. Cellular [BJS97, WN94, AAN96, GUP98, NHK98, RFS96, TSS95, ZN99, Gar99]. cellular/PCS [Gar99]. Center [HHP91, GHK^{+}93, LB97, WN94]. center-to-base [LB97]. centers [HJKN95]. centric [DDHQ98, MMO^{+}98]. century [GLDT98, Mid97]. certification [HK95]. challenge [HS95]. Challenges [RT90, ELMS99, JALR95, JPT95, MJSS95, SMS95]. Changes [JAC91]. Channel [AAN96, GCJM93, SSW93]. channels [WPE96]. character [JSS^{+}95]. characterization [EFG^{+}97]. Characterizing [SS91, CGL^{+}94]. China [WCW99]. Chip [GLDT98, AAE^{+}93]. Chip-scale [GLDT98]. CIF [ZYB98]. CIF-to-QCIF [ZYB98]. CIM [Nis99]. CIM/DEN [Nis99]. Circuit [AHM^{+}92, Ell92, ABD^{+}90, LCL99, MSHT97]. Circuit-Board [AHM^{+}92]. Circuits [CA91, FKRW92, KK92, WYKO93]. Ciruit [HCWW91]. Class [BGH^{+}93, GUP98, YLB96]. Classic [BMW94]. cleaning [WDM^{+}92]. coax [DDM^{+}96, GP96a, GMSW96]. code [BDK96]. codec [AAE^{+}93]. coded [WPE96]. coders [CKC^{+}95]. Coding [GCJM93, JLP90, ACD^{+}93, Yu98]. cognitive [JW96]. collaboration [AHRS95]. combined [BMHS98]. commerce [BBE^{+}99]. Communication [EA97, BDK96, CFH^{+}99, Fos96, GS99, LK95]. Communications [ANP93, BF90, EFG^{+}97, ALS99, Bau99a, BKcL^{+}99, CS96, CCS95, KR95, LB97, Net97, SC99, Sip95, SH99]. comparison [DDHQ98]. COMPAS [BC90]. Competencies [Pat90]. Competing [GLM92]. Competitive [FF92, GF90, MSSU96]. Competitiveness [NA91]. complexity [DER97]. compliance [HMC90]. Components [HKBW96, Kyt96]. compression [JC95]. Computer [FSV91, SKS91]. Computer-Aided [SKS91]. Computing [BH91, DM95, HJ97, Net97, Xu97]. Concepts [Hen90, DK90]. Concurrent [GLH^{+}94]. conferencing [BMHS98, HLMS93]. configuration [Mid97]. Conformance [BRW90, DS90]. Connection [HW98, Vee97]. connects [JPMK99]. conscious [WS91]. considerations [FKK93, Rab95]. considering [DCGS95]. Constructing [HE98]. Construction [Tul92]. Consumer [SS92]. consumers [SW95]. content [CH96]. continuous [BSP97]. contract [VNB91]. control [BCEN99, BHZ97, HW98, SK96, Vee97]. controls [DHVS^{+}98]. converged [BCH^{+}99, FKLS98]. convergence [BBE^{+}97, LR99]. conversion [ZYB98]. copper [LSWZ96]. Core [Pat90, HW98]. Corporate [RRW91, RHNS91, EF96]. Correction [SSW93]. cost [ESWW96, Ufo99]. coupled [Nag98]. craft [HHMP94]. creating [BLLR94, PBBV99]. creation [ABB^{+}97, Bau99b, MCG^{+}91]. critical [LMN90]. cross [JPMK99]. cryptography [Odl94]. Current [ALZ96, SAES95]. customer [BCG^{+}94, Bro91, BBH^{+}99, HHJ^{+}94]. Customers [WGHL93, ALZ96]. Customized [YS95]. cyber [CJST99].

Dalí [BRSS97]. data [ABD^{+}98, BCH^{+}99, BJS97, BSP97, CCS95, DDM^{+}96, DDHV^{+}99, EF96, EKK^{+}97, GP96b, GMSW96, HJ97, KBB99, KBP97, NHK98, RMR^{+}97, TCG^{+}98, WWS^{+}99]. Database [Sch98, Whe96]. datalinks [Was92]. DCT [ZYB98]. Decade [ANP93]. Decision [DFLN92]. Definition [GWB92, Gel94]. DEFINITY [HJKN95]. DEFINITY(R) [CLN99, SH99]. delivering [WWS^{+}99]. demand [AHK^{+}93]. Denver [CFJ^{+}90, KDF90]. Deployable [JPT95]. deployment [Bau99b, BGM^{+}95, LMW95]. Design [BGH^{+}93, BDNT96, BJL92, CFW92, FFM93, GWB92, HMC90, HPR97, IPU91, JAC91, LMW95, NL93, SAES95, SKS91, WMS95, AL95, AKM94, BCG^{+}94, CC96, CDT98, DK93b, DDH^{+}99, DER97, FHMOS97, FKK93, GCS95, HW98, JW96, MMR98, PGGA^{+}92, WMGL98, vdBC97, GF90, Nag98]. Design-by-Use [GWB92]. DESIGNER [MMR98]. Designing [Bal93, BDH^{+}98, CLRW95, DCGS95, GUP98, MC91, OT93]. designs [EP97]. desktop [BMHS98, TWN95]. destructive [CGL^{+}94]. detection [EET95, HMH99]. Developing [LMN^{+}92, SF91]. Development [BJL92, FHM^{+}91, FE91, FFM93, GLM92, HMMS90, Tuc93, AG98, BC90, BCW90, BKSZ90, BFF^{+}91, CLN99, CC96, FGJL97, GH97, MSSU96, Nag98, Pee91, YBH97]. development-process [BC90]. Device [MAB^{+}97, CDK^{+}94]. devices [CGL^{+}94]. DFX [GF90]. dial [CHV99]. dial-up [CHV99]. differentiated [BV98, Wei98]. Digital [BKR^{+}93, CB95, GRL^{+}91, KBB99, KRS98, RMR^{+}97, SD90, BBR^{+}96, BM91, BJS97, CDK^{+}94, NHK98, WN94]. digital-mobile-services [WN94]. Dimensional [AHM^{+}92, AsK93]. dimensions [JALR95]. Directions [HR93, Wal90, CS96, KR95, SK91]. directory [Nis99]. directory-hosted [Nis99]. DirSA [KK98]. Distance [ABI^{+}93a]. distances [LMMT95]. Distributed [PSS96, Bau99a, CBH96, Kyt96, PR99, vSV97]. Distribution [HE93, Max94, CCJ^{+}99]. diverse [BCH^{+}99]. Diversity [SSW93]. Dividing [CFN96]. Document [Max94, AsK93]. documentation [Bal93]. domain [AG98, ZYB98]. Doped [ZGSD92]. down [ZYB98]. down-conversion [ZYB98]. driven [HMH99]. drop [GS99]. Dynamic [ABD^{+}98].

Earth [LMMT95]. ease [Far93, OT93]. easy [Bal93]. easy-to-use [Bal93]. echo [BMHS98]. echoing [Bro91]. Eco [Nag98]. eco-design [Nag98]. Eco-supplier [Nag98]. ecology [AL95]. economic [DM95]. edge [CCS95, LMN90]. effects [LTM97]. Efficiency [FKRW92]. Efficiently [CLW95]. EGPRS [BCEN99]. Electronic [Max94, BBE^{+}99, DMW94, ESWW96, YBC^{+}94]. Electronics [ML90, BP97, Net97]. Element [HCWW91, Fos96, May92, WJC99]. Elements [BCM91, JN99]. EMC [HMC90]. Emerging [Cro94, ELMS99]. empirical [YBH97]. empowered [GS92]. emulation [CT97b]. enabler [GLH^{+}94]. enabling [BMW94]. encapsulation [BBD^{+}97]. end [BKSZ90, Far92]. Energy [OOM95]. engine [PBBV99]. Engineering [BDRS99, DM95, AKM94, AG98, BBCLS97, GLH^{+}94, HPR97, HHMP94, JW96, Nag98, OT93]. engineering-designing [OT93]. enhanced [ABL^{+}98, BDGH95, MJP^{+}96]. enhancements [CWH^{+}96]. ensuring [BFF^{+}91]. enterprise [Bau99a, SH99]. Environment [Ell92, SKS91, Sch98, AL95, BV98, Cop91, Fos96, HJ97, KHD90, MSSU96, OOM95, RW98, Sei95, SEKS94, TB98, WMS95, YBH97]. Environmental [GG90, HK95, Nag98, PH92, BD95, LTM97]. Environmentally [WS91]. environments [BCW90, DM95, Xu97]. Equalization [SSW93]. Equipment [IPU91, AFH90, BEF^{+}95]. Erbium [ZGSD92]. Erbium-Doped [ZGSD92]. Errata [Ano98e]. Erratum [Ano97e]. Error [SSW93]. errors [Jac98]. Estimating [CH96]. etching [LCL99]. E^{T M} [WJC99]. evaluation [JW96]. event [BBKS98]. event-processing [BBKS98]. Evolution [BCW90, CT97a, FSV91, FKCR92, MASY91, RHNS91, Scr94, SBH^{+}93, WBE^{+}91, ACFM94, BLR95, Cop91, HS95, KMS98, KBP97, WCW99]. Evolutions [BCM91]. evolving [SS91]. exceeding [HHJ^{+}94]. Excellence [CFJ^{+}90, CS90]. eXchange [EFG^{+}97, Bau99a]. Execution [Hen90, LRW90, ABD^{+}90, AFH90, BGM^{+}95, BCH^{+}90, DK90]. exhaustive [Wu91]. expanding [GZ96]. expectations [HHJ^{+}94]. Experiences [CBH96, CLG92]. experiment [BCF^{+}96]. Experimental [BF90]. experiments [Chr99]. extrapolation [LCK^{+}98].

Facilitating [FSW98]. facsimile [Teb99]. Factors [WGHL93]. fading [Fos96]. Fast [DFLN92, ANN97, GLH^{+}94]. fault [CH96, HKBW96, Yu98]. faults [YBH97]. Fiber [BKNP92, ZGSD92, BBH^{+}99, DDM^{+}96, GP96a, GMSW96, PGGA^{+}92, SSZS99, WCW99]. fiber-coax [DDM^{+}96, GP96a, GMSW96]. fiber-to-the-customer [BBH^{+}99]. Fibers [Wal90, Ref99]. FiberVista [WWS^{+}99]. fifty [Pen97]. Filtering [SD90]. filters [Jac98]. Final [LRW90]. Final-Assembly [LRW90]. finite [FHMOS97]. finite-state [FHMOS97]. fix [CH96]. fix-on-fix [CH96]. flawless [BGM^{+}95]. Foundation [FF92, Ros97]. frame [RFS96]. framework [BBE^{+}99, RM97, vdBC97]. frameworks [CHKP99]. France [ELC92]. Free [HM92]. Free-space [HM92]. front [BKSZ90, Far92]. FT [DJ92, YS95]. FT-2 [YS95]. FT-2000 [DJ92]. FTTC [WWS^{+}99]. FTTH [WWS^{+}99]. Future [KR95, ALS99, BP97, DER97].

gap [SS98]. gateway

[KRZ98, Teb99, TSS95, WFK^{+}98, WJC99]. general [Dem99]. generating [KMS98]. Generation [HBCS93, ZGSD92, ABDH99, Bau99b, BCC^{+}95, CHKP99, DS90, DDHQ98, DGM98, EET95, FJL^{+}96, KLR98, KMS98, KJPR97, PBBV99, Pit99, RZ99, SAES95, TB98, WMGL98]. Generator [Tot91]. Germany [HK95]. giga [WYKO93]. giga-scale [WYKO93]. gigabit [EP97]. GIS [HK95]. Global [FFM93, HE93, LMN^{+}92, SBH^{+}93, WBE^{+}91, BD95, DK93b, GH97, KNS^{+}92, ZMRR95]. Globalization [SV92, May92]. GlobeView(R) [HW98]. GlobeView(R)-2000 [HW98]. goals [BD95]. government [VNB91]. grades [CLW95]. Graphical [CBA93]. Green [DDS^{+}95, GCS95]. group [ANN97]. growth [KD97]. GSM [Dem99, RFS96]. guarantees [BDH^{+}98].

Hand [Ell92]. handoffs [GS92]. handovers [AAN96]. Happened [Tow99]. Hardened [JY91]. Hardware [BGH^{+}93, BFF^{+}91, Cop91]. harmonization [Pit99]. heart [GHJ98]. heart-beat [GHJ98]. heterogeneous [BBB^{+}99, SEKS94]. hidden [AsK93]. Hierarchical [MZ97]. High [Chr99, Ell92, KDF90, RMR^{+}97, SD90, ABD^{+}98, BDK96, CCS95, DDHV^{+}99, EKK^{+}97, GLH^{+}94, GLDT98, KM97, LTM97, RW98]. High-capacity [Chr99]. high-level [GLDT98]. High-Mix [Ell92]. High-Option [KDF90]. high-performance [KM97]. high-quality [GLH^{+}94]. high-rate [BDK96]. High-Speed [SD90, RMR^{+}97, ABD^{+}98, CCS95, DDHV^{+}99, EKK^{+}97]. high-tech [RW98]. history [BFW97]. home [LSWZ96, TB98]. hosted [Nis99]. Human [BF90, WGHL93, YLB96]. Human-Factors [WGHL93]. Human-Machine [BF90]. Humanet [BF90]. Hybrid [GP96a, BCRV95, DDM^{+}96, GMSW96].

IC [DER97]. idea [CKC^{+}95]. identification [BM91]. IEEE [SK96]. II [BLWW94, KM97]. Image [ACD^{+}93]. images [WPE96]. Imaging [AHM^{+}92]. Impact [GHN^{+}99, GG90, DBL96, GUP98, Net97]. impact-tolerant [GUP98]. impacts [Pen97]. implement [CFH^{+}99]. implementation [FHMOS97, Nis99]. implementing [Wu91]. implications [WMGL98]. importance [AL95]. Improve [NA91, OT93]. Improved [HMD^{+}95]. Improvement [Elm92, PH92, Gel94]. Improving [BKSZ90, DMM^{+}95, FKRW92, RW98, SS92]. in-building [CS98]. Incremental [SKS91, BCEN99]. independence [CCJ^{+}99]. Indoor [GCJM93]. industrial [AL95]. industry [HJ97]. Inferno^{T M} [DPP^{+}97]. Information [CC96, KDF90, AEK^{+}96, BBC^{+}90, GP96b, JW96, LPS98, Nis99, YLB96]. infrastructure [AEK^{+}96, GHN^{+}99]. Initiation [SR98]. Injection [Tul92]. innovation [Ano96a, Ano96b, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d]. Innovative [Ber93]. Installation [KGC95]. Integrated [ABB^{+}97, FKRW92, Cop91, GHN^{+}99, KK92, MSHT97, RM97, WYKO93, YLB96]. Integrating [Whe96]. Integration [DK90, LRW90, AKM94, GLDT98, SH99]. intelligence [BLR95]. Intelligent [BCM91, BKcL^{+}99, EMS91, GRL^{+}91, HCWW91, MASY91, RAMS^{+}91, SK91, SV92, WKS^{+}91, WBE^{+}91, BDGH95, BBB^{+}99, CK94, FGJL97, HZ97, KNS^{+}92, KMS98, MCG^{+}91]. Interactive [ASSS96, BM91, FKF90, SW95, HE98, LMMN95]. interchange [ABL^{+}98]. interconnect [BF97]. Interconnection [BGT90, NL93]. Interface [FM95, FSV91, BCRV95]. Interfaces [HR93, Mil93, BB93, CBA93, HE98, OT93]. International [FKK93, HCS93, LMN^{+}92, WKS^{+}91, FGJL97, FKCR92, May92]. Internationalizing [Sch95]. Internet [HMH99, RM97, SR98]. internetworking [CT97a]. interoperability [BCRV95]. interpretive [KHD90]. Interspan [CCS95]. Interval [HS98]. interworking [KRZ98]. intranets [RM97]. Introducing [LPH97, HZ97]. Introduction [AF92, Fos91, Mit94, AG98, ALS99, KK98, LCL99]. Intuity [FM95]. Intuity^{T M} [ABL^{+}98]. Inventory [BDNT96]. investment [Ber97]. involving [BCF^{+}96]. IP [AMRMV99, Bau99a, BDH^{+}98, BMB^{+}99, CDT98, DDHQ98, FKLS98, SS98, SW99, Teb99, WMGL98]. IP-centric [DDHQ98]. IRAPI [FM95]. IS-95 [CWH^{+}96]. ISDN [ABI^{+}93a, BCRV95, BF90, CHJ93, DK93b, LMN^{+}92, ULS90]. ISHMAEL [Cop91]. ISM [BDK96, Tuc93]. Issues [BLLR94, FO95, GG90, WGHL93, DM95]. iterative [Gel94].

JAM* [YS95]. JPEG [WPE96].

JPEG-coded [WPE96].

Key [CLS90, MSHT97, SBH^{+}93, ABDH99, BCRV95, GF90, Odl94]. Key-System [SBH^{+}93]. knowledge [BMW94]. knowledge-based [BMW94].

label [ESWW96]. Laboratory

[WMRG90, Scr94]. Labs

[Ano96a, Ano96b, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Spr98].

LAN [CT97b, KM97, Tuc93]. Large [GLM92, LGPR93, BDR^{+}92, BKSZ90, Ber96, BDRS99, CJST99, CP92, LMMT95, YBH97]. Large-Scale [GLM92, BDR^{+}92, BKSZ90, LMMT95]. Large-Vocabulary [LGPR93]. Laser [Tow99, KRS98]. lasers [GHK^{+}92]. layer [BBD^{+}97, Pee91, SK96, ULS90]. Layered [Fos96]. Layering [RMN97]. leading [BGM^{+}95, CCS95, LMN90]. leading-edge [LMN90]. Learning [ABI^{+}93a, RW98]. length [DDM^{+}96]. level [BMB^{+}99, GLDT98, YBH97]. LIBRA [RM97]. Lightwave [CKL92, DJ92, ZGSD92, Chr99, GS99, McG92, Run92]. Like [DMM^{+}95]. likely [Pen97]. limits [BKP90]. line [Nag98]. link [BCEN99, DDHV^{+}99]. lithographic [RT90]. lithography [SF91]. LLC [Dem99]. LLC-MAC [Dem99]. load [HSGR99]. loading [Kha97]. local [GMSW96, RZ99]. Logic [AHR91]. Loop [BKNP92, KBB99]. Lot [Ell92]. Low [TCG^{+}98, ESWW96]. low-cost [ESWW96]. Low-speed [TCG^{+}98]. Lucent [JN99, LR99].

MAC [Dem99]. Machine [ABS92, BF90, BJL92, FHMOS97]. Machines [CLRW95]. macro [Mit94]. main [BRSS97]. maintenance [Cop91, KGC95, LK95]. MAKCAL [Tot91]. making [MJP^{+}96]. Management [BBB^{+}99, DG99, FE92, GS92, AAN96, BCH^{+}99, Bar94, BDPU94, BLWW94, BGK^{+}97, BDRS99, CT97b, CHKP99, CP92, CBA93, ELMS99, FKCR92, Gar99, GHN^{+}99, HSGR99, MR91, Mid97, Nis99, OOM95, Rex91, SC99, SW99, WJC99]. manager [BRSS97, ABDH99, LMMN95]. Managing [CJST99, DG99, DER97, Elm92, JAC91, SC99, Sie91, VNB91, DM95, GH97, KD97]. Mandatory [HCS93]. Manufacturing [ABD^{+}90, AFH90, BCH^{+}90, CFJ^{+}90, CS90, FKRW92, FO95, Hen90, KDF90, Tul92, DDS^{+}95, EMS91, WS91]. Market [FFM93, HE93, IPU91, SS92, AHK^{+}93, Cro94, Far92]. markets [DK93b]. Markov [AsK93]. MARS [AHR91]. Massively [BH91]. match [HMH99]. Material [Kin90]. Materials [CG90, CLS90, GG90, Pat90, BGT90, CGL^{+}94, DCGS95, RT90]. measurements [YSH90]. mechanical [BHZ97]. mechanisms [BLLR94]. MediaBuilder [Bau99b]. MediaXact^{T M} [BBE^{+}99]. medium [SK96]. Meeting [BD95]. Memory [SD90, BRSS97, EP97]. Memory-Based [SD90]. message [ABL^{+}98, BCD^{+}94]. messaging [ABL^{+}98, Cro94, DMW94, PBBV99]. method [ANN97]. methodologies [BRW90, CDK^{+}94]. methodology [BFF^{+}91]. Methods [CA91, CLG92, HZ97]. Metro [CFR^{+}99]. Microelectronics [JY91, Cle97, HJ97, RMN97]. microprocessor [BFW97]. MIDAS [BEFK91]. middleware [EA97]. mistrust [BLLR94]. Mix [Ell92, HMH99]. mix-and-match [HMH99]. Mixed [Sub91]. Mixed-Mode [Sub91]. MMCX [EFG^{+}97]. Mobile [AEK^{+}96, CHV99, BKR^{+}93, GHK^{+}93, KJPR97, LB97, PSS96, TB98, WN94, Xu97]. Mode [CCS95, Sub91]. model [CH96, Gel94, Nis99, RW98, Ufo99]. modeling [GH97]. models [AsK93, LPS98]. Modular [MR91]. Modulation [SSW93, BDK96]. Modules [Zim93, GLDT98]. Mold [Tul92]. Molded [Zim93]. MONET [JN99]. monitor [GHJ98]. monitoring [HMH99, RZ99]. Mount [BJL92]. mousetrap [MG91]. moving [Rex91, YLB96]. multi [DG99, EA97, Fos96]. multi-element [Fos96]. multi-party [EA97]. multi-vendor [DG99]. multicasting [BSP97]. Multichip [Zim93]. Multicomputer [AHR91]. Multilayer [Jac98]. Multilingual [Spr98].

Multimedia [AHRS95, ABI^{+}93a, BB93, BE95, JALR95, Teg95, AAE^{+}93, Bau99b, BDGH95, BCC^{+}95, CWH^{+}96, EA97, HLMS93, SW95, vdBC97, EFG^{+}97, ABL^{+}98]. multinational [HMC90]. multiple [CLW95, DDM^{+}96, SSM94]. multiplexer [GS99]. multiplexing [BBD^{+}97]. multiservice [BV98, CDT98, MMR98].

National [NA91]. Net [MASY91]. Net-2000 [MASY91]. NetCare(R) [DG99]. Network [BCM91, BF90, DG99, ELMS99, HCWW91, LPS98, LK95, MASY91, SEKS94, SV92, VPR^{+}90, WJC99, WKS^{+}91, WBE^{+}91, ABDH99, ANN97, ASSS96, BCH^{+}99, BV98, BDGH95, BLWW94, Ber93, BLR95, BF97, CDT98, CCJ^{+}99, CK94, CBA93, DDH^{+}99, EKD93, FGJL97, FKLS98, GP96b, GHN^{+}99, HE98, HZ97, HSGR99, JN99, KNS^{+}92, KMS98, KJPR97, LMMN95, MCG^{+}91, Nis99, PBBV99, RAMS^{+}91, SK91, SC99, SSM94, ULS90, WMGL98, YFF94, ZN99]. Networking [FSV91, MC91, RRW91, ASFG^{+}98, ABN^{+}99, ABL^{+}98, AMRMV99, CFR^{+}99, JPMK99, JN99, MZ97, Ref99, RHNS91, ZMRR95, vdBC97]. Networks [GRL^{+}91, AAN96, ABD^{+}98, BCH^{+}99, BBR^{+}96, BDH^{+}98, BBB^{+}99, BMB^{+}99, BJS97, BDRS99, Bus98, CJST99, CDT98, CHKP99, DG99, DDM^{+}96, DHVS^{+}98, DDHV^{+}99, DGM98, FKCR92, FJL^{+}96, FLJL^{+}98, GP96a, Gar99, GS99, GMSW96, KD97, KBB99, KRZ98, KMS98, MMO^{+}98, MMR98, NHK98, Noh98, PSS96, RZ99, SS91, SW99, SSZS99, Vee97, WCW99, WFK^{+}98]. Next [DGM98, FJL^{+}96, HBCS93, ZGSD92, ABDH99, Bau99b, DDHQ98, EET95, KMS98, PBBV99, Pen97]. Next-generation [DGM98, ABDH99, Bau99b, DDHQ98, KMS98, PBBV99]. Nineties [FSV91, MC91, WS91]. NMS [BLWW94]. Node [HCWW91]. noise [MAB^{+}97]. Non [CGL^{+}94, SMS95]. Non-destructive [CGL^{+}94]. non-repeatered [SMS95]. North [GS97]. notation [Mit94]. Novel [VPR^{+}90].

OAM&P [BCM91]. OATS [BPP92]. Object [CFW92, KSS93, BDNT96]. Object-Oriented [CFW92, KSS93, BDNT96]. objective [HMH99]. objective-driven [HMH99]. objects [CBH96, vSV97]. OC [DJ92]. OC-48 [DJ92]. oceanic [LMMT95]. On-site [KBP97]. OneVision(R) [LPS98]. online [Bal93]. Only [DMM^{+}95]. operating [DPP^{+}97]. operation [LK95]. Operations [FF92, LRW90, Sta94, BCF^{+}96, CBA93, DK90, KHD90, Scr94]. opportunities [KMS98]. Optical [ASFG^{+}98, DDH^{+}99, JSS^{+}95, JPMK99, NL93, PGGA^{+}92, Ref99, SSZS99, Wal90, Was92, ABN^{+}99, AMRMV99, CGL^{+}94, CFR^{+}99, Jac98, JN99, SC99, WCW99]. optically [LMW95, MJSS95]. Option [KDF90]. organizational [GS92]. organizations [HC96]. organizing [BGK^{+}97]. Oriented [CFW92, KSS93, BDNT96]. OSI [BRW90]. other [GUP98]. outages [Ber93]. overloads [CLW95]. Overview [BG97, FE92, GZ97, Mar97, TZ99, CMS93, EFG^{+}97, FE91, FS96, LMMN95].

Packaging [ML90]. Packet [AM97,

BBD^{+}97, BF97, BJS97, Dem99, DHVS^{+}98, DDHV^{+}99, FSW98, FLJL^{+}98, LB97]. packet-based [LB97]. PacketStar^{T M} [FKLS98, WFK^{+}98]. Packs [Ell92, ABD^{+}90]. paradigm [DM95, FHMOS97]. Parallel [BH91]. part [DCGS95]. partitioning [MZ97]. Partnering [May92, ELC92]. party [EA97]. past [DER97]. path [ANN97]. PathStar^{T M} [FSW98]. Patterns [HC96, vdBC97]. PBX [FKK93]. PCS [ASSS96, GS97, Gar99, NHK98, WN94]. Pen [Mil93]. Pen-Based [Mil93]. Performance [DDJ^{+}99, FF92, BDH^{+}98, DHVS^{+}98, EFG^{+}97, FLJL^{+}98, HMH99, KM97, KD97, LB97, SS91]. Personal [CHJ93, YBH97]. personalized [BBE^{+}99]. Perspective [NL93, BLR95, FD90, Sta94]. PEST [Wu91]. phone [TCG^{+}98]. phones [GUP98]. Photonic [CG90, KK92]. photonics [HM92]. pie [CFN96]. Place [BJL92]. Planning [Kin90, LRW90, Sch98, DK90, LK95]. plans [Rex91]. plasma [LCL99]. plastics [DCGS95]. platform [Bau99b, BCC^{+}95, FLJL^{+}98]. Platforms [SBH^{+}93, WKS^{+}91, BCF^{+}96, BE95, RAMS^{+}91]. platforms-provisioning [BCF^{+}96]. PMX [BPP92]. PMX/StarMAIL [BPP92]. policies [AAN96]. Policy [HSGR99, SW99]. Policy-based [HSGR99, SW99]. Polymers [ML90]. polynomial [Kha97]. portable [GUP98]. position [BDK96]. potentials [LMMT95]. practical [ABN^{+}99]. practices [CC96, LK95]. Precursors [MVC90]. Premises [HE93, KRZ98]. present [DER97]. preventable [YBH97]. preventing [Ber93]. prevention [Yu98]. price [ESWW96]. prioritization [TB98]. private [Bus98, CHV99, MZ97, MMR98]. proactive [FJ94]. problem [Kha97]. procedure [BFF^{+}91]. Process [AF92, DFLN92, GWB92, BC90, BKSZ90, BBCLS97, FHM^{+}91, Far92, Gel94, Sie91]. Processes [CLS90, FE92, GG90, BGT90, FE91, RT90, RMN97, WDM^{+}92]. Processing [CG90, MVC90, Pat90, VPR^{+}90, BBKS98, BKP90, Che98, CK94, FD90, LMN90, PLS90, PSS96, Rab95]. Processor [FO95, BGM^{+}95]. Processors [CB95]. Product [AF92, GWB92, HCS93, HMD^{+}95, IPU91, CKC^{+}95, DDS^{+}95, GLH^{+}94, GCS95, JW96, LMN90]. production [HHMP94]. productive [HC96]. Productivity [Tul92, RW98, YSH90]. Products [SS92, VPR^{+}90, BB93, GF90, GUP98, MJSS95, SMS95, YFF94]. professionals [YLB96]. profitable [GF90]. Program [BJL92, CFH^{+}99]. Program-and-Place [BJL92]. Programming [CFW92, FM95]. Progressive [Ell92]. Project [FE92, YS95, CP92, KRS98, MR91, Rex91, Sie91]. projection [SF91]. projects [Sch95, Sul96]. ProLogix [CLN99]. promise [Kyt96, LR99]. proposal [BCEN99]. prose [BMW94]. prospectus [DK93a]. protection [ANN97, CLW95]. protocol [CT97b, DS90, DDM^{+}96, DDHV^{+}99, Hol90, ULS90, LMMN95, SR98, Teb99]. Protocols [AMRMV99, DHVS^{+}98, BRW90, Xu97]. Prototyping [YS95]. provider [WFK^{+}98]. Providing [HCWW91, CLW95, CCJ^{+}99, SR98]. provision [BDGH95]. provisioning [BCF^{+}96]. Pseudo [AsK93, Wu91]. pseudo-exhaustive [Wu91]. Public [Odl94, EKD93]. PWT [TCG^{+}98].

QCIF [ZYB98]. QFD [Bro91]. QoS [BV98, Bus98, FLJL^{+}98, MK98, Wei98]. QoS-capable [Bus98]. Quality [Elm92, FE92, HMD^{+}95, PH92, Tul92, BDPU94, BFF^{+}91, CP92, FJ94, GLH^{+}94, RW98]. Quantum [GHK^{+}92]. queries [HZ97]. Quick [YS95].

R [WJC99]. R&D [VNB91]. R/E^{T M}

[WJC99]. Radiation [JY91].

Radiation-Hardened [JY91]. Radio [SD90, BCEN99, BKR^{+}93, Dem99, ESWW96, KLR98]. random [Jac98]. rapid [BCC^{+}95, CLN99]. rate [BCRV95, BDK96]. ray [SF91]. Reaching [BKP90]. Real [HMH99, SK96, Teb99, WMRG90, BBKS98]. Real-time [HMH99, SK96, Teb99, BBKS98]. reality [MJP^{+}96, Teg95]. realization [GLH^{+}94]. Realizing [LR99]. Reaping [KSS93]. Recognition [LGPR93, WGHL93, WMRG90, AsK93, JSS^{+}95, JPT95]. recyclability [DCGS95]. recycled [DCGS95]. Redefining [Bau99a]. reduction [HS98]. redundancy [BCEN99]. Reengineering [AF92]. rejuvenation [HKBW96]. relation [FGJL97]. relay [BLWW94]. Reliability [CLS90]. Reliable [BSP97, Xu97]. repeatered [MJSS95, SMS95]. replacement [Mit94]. requirements [HPR97, HS98, SAES95, Teb99]. Research [BSW92, BE95, Pee91]. Residential [GMSW96]. Resources [Pat90, CLW95, YLB96]. Response [FM95, HRRS90]. restoration [DDH^{+}99]. retrofit [KBP97]. Reusable [CHKP99, BDNT96]. reuse [BDR^{+}92, LPH97]. revenue [KMS98]. reviews [FHM^{+}91]. Revolutionizing [HJKN95]. RF [MAB^{+}97]. ring [Kha97]. Robust [BPP92, CLS90, GCJM93, BCG^{+}94]. role [KMS98, LMN90]. root [Yu98]. routing [FKCR92, FLJL^{+}98, RM97].

Sandia [NA91, Pee91]. satisfaction [BCG^{+}94]. Scalable [Che98]. Scale [GLM92, BDR^{+}92, BKSZ90, FSW98, GLDT98, LMMT95, WYKO93]. scan [YBC^{+}94]. scenarios [Noh98]. scheme [Kha97]. science [FD90]. Scientist [Tow99]. Scoping [Sul96]. SDH [BDRS99, Noh98]. SDH/SONET [Noh98]. Secure [SSM94, AKM94]. Security [BK94, DMW94, MG91, CJST99, SEKS94]. self [BGK^{+}97, JSS^{+}95, JW96, Wu91]. self-organizing [BGK^{+}97]. self-service [JSS^{+}95, JW96]. self-test [Wu91]. semi [WDM^{+}92]. semi-aqueous [WDM^{+}92]. Semiconductor [MVC90, CGL^{+}94, Pee91]. Sequencing [LRW90]. Sequential [CA91]. server [ABL^{+}98, FSW98, SH99]. Service [CK94, FKRW92, HCWW91, MMO^{+}98, MCG^{+}91, ABB^{+}97, Bau99b, BCD^{+}94, CLW95, Dem99, JSS^{+}95, JW96, MJP^{+}96, RM97, WFK^{+}98, BDNT96, CCS95, MASY91]. service-based [RM97]. Service-centric [MMO^{+}98]. Services [HCWW91, IPU91, SV92, ASSS96, ALS99, BV98, BBE^{+}99, BDGH95, BCF^{+}96, BB93, CWH^{+}96, CHV99, CCS95, CS98, DK93a, DDM^{+}96, FCT^{+}99, GMSW96, HZ97, KMS98, PBBV99, RMR^{+}97, SR98, SW95, Wei98, WN94, WWS^{+}99, DG99]. Serving [WGHL93]. Session [SR98]. set [AAE^{+}93]. shared [CLW95, EP97]. shared-memory [EP97]. shift [DM95]. Signal [CB95, BCF^{+}96]. signatures [BM91]. Silicon [Cle97, BKP90, MAB^{+}97, Ros97, RMN97, WYKO93]. Silver [Gel94, DBL96]. simple [DDHV^{+}99]. Simulation [AHR91, BEFK91, BGK^{+}97, FLJL^{+}98]. site [KBP97]. Small [Ell92, MSSU96, BBD^{+}97]. Small-Lot [Ell92]. Small-team [MSSU96]. SMART [KD97, SSM94]. soft [SF91]. Softswitch [LR99]. Software [BH91, Ber96, Ber97, FHO92, GLM92, HHMP94, Mid97, Rex91, YSH90, AG98, BCW90, BBC^{+}90, BSW92, BDNT96, BKSZ90, BBCLS97, CH96, CFN96, CP92, Cop91, FJ94, GH97, GHJ98, HC96, HKBW96, KK98, LPH97, Sch95, Sul96, YBH97, Yu98, HMMS90]. software-development [BKSZ90]. software/hardware [Cop91]. solid [BP97, Net97, Pen97]. solid-state [BP97, Net97, Pen97]. solution [BCH^{+}99, EKK^{+}97, KD97, WFK^{+}98]. Solutions [HHP91, MC91, CLN99, CCJ^{+}99, KBB99]. Some [Pen97]. SONET [BDRS99, Kha97, Noh98]. SONET/SDH [BDRS99]. space [Fos96, HM92, LTM97]. space-based [LTM97]. space-time [Fos96]. spatial [HJ97]. spectrum [BGK^{+}97]. Speech [CKC^{+}95, FD90, FCT^{+}99, GCJM93, LGPR93, VPR^{+}90, WGHL93, WMRG90, JLP90, JPT95, RFS96, SO95, Spr98].

Speech-based [FCT^{+}99]. Speed [SD90, ABD^{+}98, CCS95, DDHV^{+}99, EKK^{+}97, RMR^{+}97, TCG^{+}98]. SQA [FJ94]. st [GLDT98]. Standards [HCS93, ACD^{+}93, FGJL97, HLMS93, Pit99]. standards-based [HLMS93]. StarKeeper [BLWW94]. StarMAIL [BPP92]. state [BP97, FHMOS97, Net97, Pen97]. station [LB97]. Statistical [GH97]. step [vSV97]. steps [MSHT97]. stereo [BMHS98]. STM [DDM^{+}96]. storage [BRSS97]. storeroom [BCH^{+}90]. stories [BMW94]. Strained [Pee91]. Strained-layer [Pee91]. Strategic [Pat90, May92]. strategy [Ber97, LCK^{+}98]. stream [Xu97]. streams [BSP97]. Stress [PH92]. Striving [CFJ^{+}90, CS90]. Structured [HE93]. Studies [LMMT95]. study [CFH^{+}99, FLJL^{+}98, KK98, YBH97]. styles [BCF^{+}96]. Styx(R) [PR99]. subassemblies [YBC^{+}94]. Subword [LGPR93]. Subword-Based [LGPR93]. Successful [AG98, CFH^{+}99]. Summary [CFJ^{+}90]. Sunrise [BBKS98]. supplier [Nag98]. supply [Nag98]. supply-line [Nag98]. support [BC90, KHD90]. Supported [CLG92]. Surface [BJL92]. Surface-Mount-Assembly [BJL92]. survivability [ANN97]. Switch [FSV91, HBCS93, LMN^{+}92, RZ99, ACFM94, BLR95, BFF^{+}91, DDJ^{+}99, Far92, FKLS98, GH97, HHJ^{+}94, KBP97, TSS95, WMS95]. Switch-Computer [FSV91]. Switch-to-Computer [FSV91]. Switched [BBR^{+}96, BF97, EKD93, FKCR92, RMR^{+}97]. switches [Che98, HW98]. Switching [HBCS93, ANN97, BDPU94, EP97, GHK^{+}93, HM92, HS95, LB97, WN94, YFF94]. switching-network [YFF94]. Synchronization [GRL^{+}91]. Synchronous [Sub91]. Synthesis [Sub91, SO95, Spr98]. Synthesized [BMHS98]. Synthesizing [CA91]. System [BEFK91, HBCS93, KDF90, KJPR97, NL93, Scr94, SBH^{+}93, AKM94, BCH^{+}99, BC90, BBKS98, BBC^{+}90, BCF^{+}96, CCJ^{+}99, DPP^{+}97, ESWW96, PLS90, PGGA^{+}92, TCG^{+}98, WWS^{+}99, BBE^{+}99, BDNT96, DJ92, RMR^{+}97]. Systematic [GHJ98]. Systems [Bar94, FHO92, HRRS90, HE93, Sub91, VPR^{+}90, ZGSD92, BCRV95, Bal93, BDPU94, Bau99a, BKSZ90, BMW94, BLLR94, CMS93, CK94, DMW94, DDJ^{+}99, EMS91, GS97, GP96b, JPT95, KGC95, KHD90, LK95, LMW95, PR99, RFS96, Run92, SAES95, SC99, Sta94, YBC^{+}94, ZMRR95].

Talking [CLRW95]. TDMA [GS97, LB97]. team [MSSU96]. teams [GS92]. Teamwork [LMN^{+}92]. tech [RW98]. Techniques [FHO92, SSW93, CGL^{+}94]. technological [FS96]. Technologies [HRRS90, LR99, NA91, BMW94, Ber93, BK94, CMS93, Cro94, EMS91, MAB^{+}97, MCG^{+}91, SMS95]. Technology [CKL92, FKF90, KSS93, NL93, Zim93, BEF^{+}95, CC96, Cle97, Far93, GZ96, JC95, JALR95, JW96, KK98, KLR98, KGC95, LCL99, LPH97, McG92, MG91, MJSS95, Pen97, Sei95, Sip95, TWN95, WS91, Whe96, YLB96]. Telecommunications [IPU91, KJPR97, BB93, GHK^{+}92, May92, TB98]. telephone [ABB^{+}97, EKD93]. telephones [KRS98]. Telephony [ABI93b, AM97, EKD93, FSW98, SR98, SS98]. telesystemes [ELC92]. TelMex [PGGA^{+}92]. Ten [OT93]. Terminal [HR93, MK98, BEF^{+}95]. test [AHK^{+}93, CDK^{+}94, Wu91]. Testable [CA91]. Testing [BPP92, FHO92, HMD^{+}95, PH92, BRW90, GHJ98, LCK^{+}98, YBC^{+}94]. tests [DS90]. Text [SO95, Spr98]. Text-to [SO95]. text-to-speech [Spr98]. their [RMN97]. Theory [Hen90, DK90]. Third [KJPR97, Pit99, TB98, CHKP99, KLR98, WMGL98]. Third-generation [KJPR97, Pit99, TB98, CHKP99, KLR98, WMGL98]. Three [AHM^{+}92, vSV97]. Three-Dimensional [AHM^{+}92]. Time [SS92, BBKS98, Fos96, HMH99, Kha97, SK96, Teb99]. TM [CLN99]. Today [CLRW95]. tolerance [HKBW96]. tolerant [GUP98]. Tomorrow [CLRW95]. Tool [Tul92, MMR98, SF91, Wu91]. Tools [BBCLS97, CLG92, FHO92, SKS91, HPR97]. topology [ABL^{+}98, CT97b]. Total [BDPU94, CP92]. Tracking [Sch98]. traffic [LB97, SK96]. training [RW98]. Transaction [HR93]. transcoding [RFS96]. Transfer [CCS95]. Transmission [BGH^{+}93, WPE96, Chr99, KLR98, SS91, Xu97]. transparent [BLLR94]. transport [ALZ96, ABDH99, ABN^{+}99, DDHQ98, SC99]. Trends [CDK^{+}94, ALZ96, ELMS99, FS96]. Trial [ABI^{+}93a]. triggers [HZ97]. Trunk [Gar99, WJC99]. trustworthy [BLLR94]. Two [ZTL96, AsK93]. two-dimensional [AsK93]. Two-way [ZTL96]. type [RM97].

UMTS [KJPR97]. Undersea [MJSS95, Run92, Sip95, SMS95, KR95, KGC95, LK95, LMW95, SAES95, ZMRR95]. underwater [EET95]. Unified [PBBV99]. unit [BDPU94]. Universal [KRZ98, KJPR97]. UNIX [Sch95]. unlicensed [KM97]. Update [CKL92]. uplink [BKcL^{+}99]. Usability [JW96, OT93]. Use [GWB92, Bal93, DCGS95, Far93, OT93]. used [BEF^{+}95]. User [BMB^{+}99, CLRW95, Mil93, BB93, OT93]. User-level [BMB^{+}99]. Using [ABI^{+}93a, BPP92, Far93, NA91, NHK98, PH92, TWN95, YS95, vdBC97, BCEN99, CH96, Fos96, GHJ98, HZ97, KRZ98, RMR^{+}97, SSM94].

V [LMMN95]. value [PBBV99]. value-creating [PBBV99]. Variability [FF92]. variable [DDM^{+}96]. VCTV [AHK^{+}93]. vendor [DG99]. verification [Hol90, ULS90]. VeriSoft [GHJ98]. very [YBH97]. via [GMSW96]. vibrations [BHZ97]. Video [ANP93, ABI93b, CHJ93, AAE^{+}93, AHK^{+}93, ACD^{+}93, BBR^{+}96, EKD93, LMMN95, ZYB98, LMMN95]. video-codec [AAE^{+}93]. video-on-demand [AHK^{+}93]. VideoPhone [EKD93]. View [Tul92]. Virtual [ANN97, Bus98, CT97b, CHV99, FHMOS97, MZ97, MMR98, TB98]. vision [ABN^{+}99, Teg95, CB95, Rab95]. visual [FLJL^{+}98, HE98]. Visualizing [EF96]. VLSI [LCL99]. Vocabulary [LGPR93]. Voice [CKLM98, CDT98, FKF90, HRRS90, KMS98, WFK^{+}98, BCH^{+}99, Bro91, DHVS^{+}98, KBB99, KRZ98, NHK98, PLS90, Rab95, SS91, WMGL98]. voice-over-IP [WMGL98]. VPN [MMR98]. vulnerabilities [CJST99].

warehouse [GP96b]. WaveLAN [Tuc93]. WaveLAN(R) [KM97]. WaveLAN(R)-II [KM97]. wavelength [GS99]. WaveStar^{T M} [ABDH99]. way [BGM^{+}95, ZTL96]. ways [OT93]. WDM [SSZS99]. web [ABB^{+}97, Whe96]. well [GHK^{+}92]. Wide [Whe96, DHVS^{+}98]. Wideband [BF90, JLP90]. WIN [FGJL97]. Wireless [BBE^{+}97, BGK^{+}97, CMS93, CS98, GCJM93, LTM97, MJP^{+}96, WMGL98, ABD^{+}98, BKcL^{+}99, CS96, CHKP99, FGJL97, Fos96, FJL^{+}96, GHK^{+}93, GZ96, HZ97, KM97, KD97, Pit99, PSS96, TCG^{+}98, Tuc93, Ufo99, WPE96, ZTL96]. wireline [BBE^{+}97]. Works [KDF90, CFJ^{+}90]. World [BGH^{+}93, WMRG90, GZ96, Whe96]. World-Class [BGH^{+}93]. worlds [vSV97].

X [GF90, SF91]. X-ray [SF91].

years [Pen97].

References

Ackland:1993:VCC

[AAE^{+}93] B. D. Ackland, R. Aghevli, I. Eldumiati, A. C. Englander, and E. Scuteri, Jr. A video-codec chip set for multimedia applications. AT&T Technical Journal, 72(1):50–66, January/February 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Agrawal:1996:CMP

[AAN96] Prathima Agrawal, Dinesh K.

Anvekar, and Balakrishnan Narendran. Channel management policies for handovers in cellular networks. Bell Labs Technical Journal, 1(2): 97–110, Summer 1996. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Atkins:1997:IWT

[ABB^{+}97] David L. Atkins, Thomas Ball, Thomas R. Baran, Michael A. Benedikt, Kenneth C. Cox, David A. Ladd, Peter A. Mataga, Carlos Puchol, J. Christopher Ramming, Kenneth G. Rehor, and Curtis Tuckey. Integrated web and telephone service creation. Bell Labs Technical Journal, 2(1):19–35, Spring 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Albano:1990:MEC

[ABD^{+}90] R. E. Albano, C. Buyukkoc, B. T. Doshi, D. X. Callaway, D. J. Friedman, T. R. Mc-Clure, L. A. Schmitt, and J. J. Svitak. Manufacturing execution: circuit packs. AT&T Technical Journal, 69(4):64–80, July/August 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Andrews:1998:DBA

[ABD^{+}98] Matthew Andrews, Simon C. Borst, Francis Dominique, Predrag R. Jelenkovic, Krishnan Kumaran, K. G. Ramakrishnan, and Philip A. Whiting. Dynamic bandwidth allocation algorithms for highspeed data wireless networks. Bell Labs Technical Journal, 3 (3):30–49, Autumn 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alegria:1999:WTB

[ABDH99] Carlos A. Alegria, Shalom S. Bergstein, Kimberly A. Dixson, and Christopher J. Hunt. The WaveStar^{T M} BandWidth Manager: The key building block in the next-generation transport network. Bell Labs Technical Journal, 4(1):42–57, Spring 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

An:1993:MDL

[ABI^{+}93a] Chungming An, Brian T. Barcelo, Joyce A. Inkrott, Arthur R. Snowdon, and Kenneth J. Trojniar. A multimedia distance learning trial using ISDN BRI. AT&T Technical Journal, 72(1):15–21, January/February 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Angiolillo:1993:VT

[ABI93b] Joel S. Angiolillo, Harry E.

Blanchard, and Edmond W. Israelski. Video telephony. AT&T Technical Journal, 72(3): 7–20, May/June 1993. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Anderl:1998:EMN

[ABL^{+}98] Ewald C. Anderl, Craig S.

Batzel, Gurmeet S. Lamba, John M. Leikness, and Bhaskar N. Sanakkayala. An enhanced message networking topology: Multimedia messaging with the Intuity^{T M} interchange server. Bell Labs Technical Journal, 3 (2):124–135, Summer 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alferness:1999:PVO

[ABN^{+}99] Rod C. Alferness, Paul A.

Bonenfant, Curtis J. Newton, Kevin A. Sparks, and Eve L. Varma. A practical vision for optical transport networking. Bell Labs Technical Journal, 4(1):3–18, Spring 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anderson:1992:ACA

[ABS92] Carl R. Anderson, Allan R.

Brown, and Gary G. Seaman. An automated cable assembly machine. AT&T Technical Journal, 71(4):24–29, July/August 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Aravind:1993:IVC

[ACD^{+}93] R. Aravind, G. L. Cash, D. L. Duttweiler, Hsueh-Ming Hang, B. G. Haskell, and A. Puri. Image and video coding standards. AT&T Technical Journal, 72(1): 67–79, January/February 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Anderson:1994:ES

[ACFM94] Thomas W. Anderson, Paul D. Carestia, James H. Foster, and Michael N. Meyers. The evolution of the 4ESS switch. AT&T Technical Journal, 73 (6):93–100, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ayanoglu:1996:MII

[AEK^{+}96] Ender Ayanoglu, Kai Y. Eng, Mark J. Karol, Zhao Liu, Pramod Pancha, Malathi Veeraraghavan, and Clark B. Woodworth. Mobile information infrastructure. Bell Labs Technical Journal, 1(2):143–164, Summer 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Arnold:1992:RNP

[AF92] George W. Arnold and Mark C. Floyd. Reengineering the new product introduction process. AT&T Technical Journal, 71 (6):12–19, November/December 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Albano:1990:MEE

[AFH90] Robert E. Albano, David J.

Friedman, and Stanley A. Hendryx. Manufacturing execution: equipment. AT&T Technical Journal, 69(4):53–63, July/August 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ardis:1998:SID

[AG98] Mark A. Ardis and Janel A.

Green. Successful introduction of domain engineering into software development. Bell Labs Technical Journal, 3(3): 10–20, Autumn 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Allen:1993:VVD

[AHK^{+}93] James R. Allen, Blaise L. Heltai, Arthur H. Koenig, Donald F. Snow, and James R. Warson. VCTV: a video-on-demand market test. AT&T Technical Journal, 72(1):7–14, January/February 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Amir:1992:TDI

[AHM^{+}92] Israel Amir, Frank P. Higgins, John B. Macdonald, Howard N. Gaston, and Rodney E. Griffin. Three-dimensional imaging for circuit-board assembly. AT&T Technical Journal, 71(4): 30–36, July/August 1992. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Agrawal:1991:LSM

[AHR91] Prathima Agrawal, Chong Hoc Hao, and Michel Remillard. Logic simulation on the MARS multicomputer. AT&T Technical Journal, 70(1):21–35, January/February 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ahuja:1995:MC

[AHRS95] Sudhir R. Ahuja, H. Fred Haisch, Ram S. Ramamurthy, and Lucinda M. Sanders. Multimedia collaboration. AT&T Technical Journal, 74(5):46–53, September/October 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Amoroso:1994:EAS

[AKM94] E. Amoroso, W. E. Kleppinger, and D. Majette. An engineering approach to secure system analysis, design, and integration. AT&T Technical Journal, 73 (5):40–51, September/October 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Allenby:1995:IIE

[AL95] Braden R. Allenby and Robert A. Laudise. The importance of industrial ecology and design for environment to AT&T. AT&T Technical Journal, 74(6):8–16, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

August:1999:IFC

[ALS99] Kit G. August, Victor B.

Lawrence, and Burton R. Saltzberg. An introduction to future communications services and access. Bell Labs Technical Journal, 4(2):3–20, Summer 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Alegria:1996:CTA

[ALZ96] Carlos A. Alegria, Hyeon J. Lee, and Richard Zoccolillo. Current trends in access and transport architectures for business customers. Bell Labs Technical Journal, 1(1):78–87, Spring 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ahuja:1997:PT

[AM97] Sudhir R. Ahuja and Krishna G.

Murti. Packet telephony. Bell Labs Technical Journal, 2(2): 5–14, Summer 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anderson:1999:PAI

[AMRMV99] Jon Anderson, James S. Manchester, Antonio Rodriguez-Moral, and Malathi Veeraraghavan. Protocols and architectures for IP optical networking. Bell Labs Technical Journal, 4 (1):105–124, Spring 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anderson:1997:VPG

[ANN97] Jon Anderson, Curtis J. Newton, and Tai H. Noh. Virtual path group protection switching — a method for fast ATM network survivability. Bell Labs Technical Journal, 2(2): 213–233, Summer 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1996:BLIa

[Ano96a] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 1(1):146–147, Spring 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1996:BLIb

[Ano96b] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 1(2):216–217, Summer 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1997:BLIa

[Ano97a] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 2(1):220–222, Spring 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1997:BLIb

[Ano97b] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 2(2):258–259, Summer 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1997:BLIc

[Ano97c] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 2(3):243–244, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1997:BLId

[Ano97d] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 2(4):169–170, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1997:E

[Ano97e] Anonymous. Erratum. Bell Labs Technical Journal, 2(3): 261, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1998:BLIa

[Ano98a] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 3(1):145–146, Spring 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1998:BLIb

[Ano98b] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 3(2):137–138, Summer 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1998:BLIc

[Ano98c] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 3(3):175–176, Autumn 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1998:BLId

[Ano98d] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 3(4):338–339, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1998:E

[Ano98e] Anonymous. Errata. Bell Labs Technical Journal, 3(1): 159, Spring 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1999:BLIa

[Ano99a] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 4(1):323, Spring 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1999:BLIb

[Ano99b] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 4(2):218–219, Summer 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1999:BLIc

[Ano99c] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 4(3):191–192, Autumn 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Anonymous:1999:BLId

[Ano99d] Anonymous. Bell Labs innovation briefs. Bell Labs Technical Journal, 4(4):273–274, October/December 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Aldermeshian:1993:VCD

[ANP93] Hrair Aldermeshian, William H.

Ninke, and Randolph J. Pilc. The video communications decade. AT&T Technical Journal, 72(1):2–6, January/ February 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Al-Salameh:1998:ON

[ASFG^{+}98] Daniel Y. Al-Salameh, Mohammad T. Fatehi, William J. Gartner, Stan Lumish, Bruce L. Nelson, and Kamal K. Raychaudhuri. Optical networking. Bell Labs Technical Journal, 3 (1):39–61, Spring 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Agazzi:1993:PTD

[AsK93] Oscar E. Agazzi and Shyh shiaw Kuo. Pseudo two-dimensional hidden Markov models for document recognition. AT&T Technical Journal, 72(5):60–72, September/October 1993. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Aprille:1996:IBS

[ASSS96] Thomas J. Aprille, Larry M.

Schwerin, James D. Sipes, and Nancy S. Stevens. Interactive broadband services and PCS network architecture. Bell Labs Technical Journal, 1(1):11–27, Spring 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Baldasare:1993:DEU

[Bal93] J. Baldasare. Designing easy-to-use online documentation systems. AT&T Technical Journal, 72(3):67–74, May/June 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Barbier:1994:SM

[Bar94] Susan Barbier. Systems management in the 1990s. AT&T Technical Journal, 73(4):82–96, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bauer:1999:IES

[Bau99a] Eric J. Bauer. IP exchange systems — Redefining distributed communications in the enterprise. Bell Labs Technical Journal, 4(2):43–56, Summer 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Baumann:1999:MNG

[Bau99b] Frank V. Baumann. Media-Builder: A next-generation multimedia service creation and deployment platform. Bell Labs Technical Journal, 4(2): 57–70, Summer 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Benimoff:1993:MUI

[BB93] N. I. Benimoff and M. J. Burns.

Multimedia user interfaces for telecommunications products and services. AT&T Technical Journal, 72(3):42–49, May/ June 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bieszczad:1999:MHN

[BBB^{+}99] Andrzej Bieszczad, Pratik K.

Biswas, Walter Buga, Manu Malek, and Hai Tan. Management of heterogeneous networks with intelligent agents. Bell Labs Technical Journal, 4(4): 109–135, October/December 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Belanger:1990:TSI

[BBC^{+}90] David G. Belanger, Ronald J.

Brachman, Yih-Farn Chen, Prekumar T. Devanbu, and Peter G. Selfridge. Toward a software information system. AT&T Technical Journal, 69(2):22–41, 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Brownlie:1997:TSP

[BBCLS97] Robert A. Brownlie, Philip E.

Brown, Kathleen Culver-Lozo, and James J. Striegel. Tools for software process engineering. Bell Labs Technical Journal, 2 (1):130–143, Spring 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Baldwin:1997:ANA

[BBD^{+}97] John H. Baldwin, Behram H. Bharucha, Bharat T. Doshi, Subrahmanyam Dravida, and Sanjiv Nanda. AAL-2 — a new ATM adaptation layer for small packet encapsulation and multiplexing. Bell Labs Technical Journal, 2(2):111–131, Summer 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bergren:1997:WWC

[BBE^{+}97] Mark H. Bergren, Brian D. Bolliger, Deborah T. Earl, Dolores L. Grossmann, Bin wen Ho, and Robin J. Thompson. Wireless and wireline convergence. Bell Labs Technical Journal, 2(3):194–206, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bartell:1999:MTS

[BBE^{+}99] Randy L. Bartell, Nadine A. Blackwood, Dominik Eggenschwiler, Mai-Huong Nguyen, Christian Schnidrig, and Michael J. Yatchman. The MediaXact^{T M} System — a framework for personalized electronic commerce services. Bell Labs Technical Journal, 4(2):153–173, Summer 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Buchholz:1999:BFA

[BBH^{+}99] D. Bruce Buchholz, Charles C.

Byers, Stephen J. Hinterlong, George D. Huensch, Rick L. Morrison, Trent R. Nanke, James P. Runyon, and Hsien-Chuen Yu. Broadband fiber access: A fiber-to-the-customer access architecture. Bell Labs Technical Journal, 4(1):282– 299, Spring 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Baulier:1998:SRT

[BBKS98] Gerald D. Baulier, Stephen M.

Blott, Henry F. Korth, and Avi Silberschatz. Sunrise: A real-time event-processing system. Bell Labs Technical Journal, 3(1):3–18, Spring 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bankapur:1996:SDV

[BBR^{+}96] Rang J. Bankapur, Hugh J.

Beuscher, James P. Runyon, Amalesh C. R. Sanku, and Chander S. Sehgal. Switched digital video access networks. Bell Labs Technical Journal, 1 (1):66–77, Spring 1996. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Barnard:1990:CDP

[BC90] H. Jack Barnard and Robert B.

Collicott. COMPAS. a developmentprocess support system. AT&T Technical Journal, 69(2):52–64, 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Beyler:1995:ABP

[BCC^{+}95] Eric R. Beyler, Olivier B. Clarisse, Edward A. Clark, Y. H. Levendel, and Robert E. Richardson. An ATM-based platform for rapid generation of multimedia applications. AT&T Technical Journal, 74(5):106– 116, September/October 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Benimoff:1994:AMS

[BCD^{+}94] N. I. Benimoff, R. E. Carlin, Jr., A. N. Daudelin, F. Remite, Jr., S. P. Shinners, and D. J. Daniro, Jr. AT&T message service. AT&T Technical Journal, 73(3):46–56, 1994. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Balachandran:1999:PER

[BCEN99] Krishna Balachandran, Keith F. Conner, Richard P. Ejzak, and Sanjiv Nanda. A proposal for EGPRS radio link control using link adaptation and incremental redundancy. Bell Labs Technical Journal, 4(3):19–36, Autumn 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Belanger:1996:ASS

[BCF^{+}96] D. G. Belanger, Yih-Farn Chen, N. R. Fildes, B. Krishnamurthy, P. H. Rank, Jr., Kiem-Phong Vo, and T. E. Walker. Architecture styles and services: an experiment involving the signal operations platformsprovisioning operations system. AT&T Technical Journal, 75(1):54–63, 1996. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Boza:1994:ACS

[BCG^{+}94] L. B. Boza, T. J. Ciaccia, D. A. Gatenby, R. W. Muise, K. K. Ng, and G. M. Yanizeski. Achieving customer satisfaction through robust design. AT&T Technical Journal, 73(1):48–58, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Burman:1990:MES

[BCH^{+}90] David Y. Burman, Ke-Tsai Edward Chin, John Huber, Gerald M. Barr, Robert D. Storrs, Kaveh Hushyar, and Barbara A. Binder. Manufacturing execution. storeroom. AT&T Technical Journal, 69(4):81–89, July/August 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bagga:1999:NMS

[BCH^{+}99] Yudhveer S. Bagga, Arati Chaudhury, Charleen E. Hird, Ted C. Lee, William W. Pretyka, Danilo L. Puseljic, Reza Shafie-Khorasani, Thomas P. Tignor, and Rulei Ting. A network management system solution for diverse and converged voice and data networks. Bell Labs Technical Journal, 4(4):

190–203, October/December 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bennett:1991:INO

[BCM91] Ronnie L. Bennett, Jerry C. Chen, and Thomas B. Morawski. Intelligent network OAM&P capabilities and evolutions for network elements. AT&T Technical Journal, 70(3–4):85–98, March/ April 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Baker:1995:IBR

[BCRV95] Albert D. Baker, James Chen, Nancy T. Radics, and Lee A. Vallone. ISDN basic rate interface interoperability with key and hybrid systems. AT&T Technical Journal, 74(4):94– 102, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Belanger:1990:ESD

[BCW90] D. G. Belanger, S. G. Chappell, and M. Wish. Evolution of software development environments. AT&T Technical Journal, 69(2):2–6, 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Borum:1995:MAG

[BD95] John C. Borum and Barry F. Dambach. Meeting AT&T’s global environmental goals. AT&T Technical Journal, 74 (6):36–43, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Beacken:1995:PIA

[BDGH95] Marc Beacken, Robert L. Doran, Michael J. Geller, and Julia Hirschberg. The provision of intelligent agent based enhanced multimedia network services. AT&T Technical Journal, 74 (5):68–77, September/October 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Benmohamed:1998:DIN

[BDH^{+}98] Lotfi M. Benmohamed, Subrahmanyam Dravida, P. Harshavardhana, Wing Cheong Lau, and Ajay K. Mittal. Designing IP networks with performance guarantees. Bell Labs Technical Journal, 3(4):273– 296, October/December 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bar-David:1996:BCP

[BDK96] Israel Bar-David and Rajeev Krishnamoorthy. Barker code position modulation for highrate communication in the ISM bands. Bell Labs Technical Journal, 1(2):21–40, Summer 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bergholm:1996:SDI

[BDNT96] J. O. Bergholm, J. M. Davis, B. Nadji, and P. D. Ting. Service Design and Inventory System — an object-oriented reusable software asset. AT&T Technical Journal, 75(2):47–57, 1996. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Basinger:1994:TQM

[BDPU94] Ronald G. Basinger, Michael J. DiMario, Thomas C. Pingel, and Michael J. Urban. Total quality management in the switching systems business unit. AT&T Technical Journal, 73(6): 7–18, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Beck:1992:ALS

[BDR^{+}92] R. P. Beck, S. R. Desai, D. R. Ryan, R. W. Tower, D. Q. Vroom, and L. M. Wood. Architectures for large-scale reuse. AT&T Technical Journal, 71 (6):34–45, November/December 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Budka:1999:ELS

[BDRS99] Kenneth C. Budka, Richard J. DiPasquale, Bruno W. A. Rijsman, and Anekal B. Sripad. Engineering large SONET/SDH management networks. Bell Labs Technical Journal, 4(4): 136–154, October/December 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Berkley:1995:MRP

[BE95] David A. Berkley and J. Robert Ensor. Multimedia research platforms. AT&T Technical Journal, 74(5):34–45, September/October 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Baumhauer:1995:ATU

[BEF^{+}95] John C. Baumhauer, Jr., Scott H. Early, John H. Fikus, Steven L. Gay, and Michael A. Zuniga. Audio technology used in AT&T’s terminal equipment. AT&T Technical Journal, 74(2):57–70, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bierbauer:1991:SSM

[BEFK91] John W. Bierbauer, Jonathan A.

Eiseman, Faiq A. Fazal, and James J. Kulikowski. System simulation with MIDAS. AT&T Technical Journal, 70(1):36–51, January/February 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bernstein:1993:ITP

[Ber93] Lawrence Bernstein. Innovative technologies for preventing network outages. AT&T Technical Journal, 72(4):4–10, July/August 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bernstein:1996:SL

[Ber96] Larry Bernstein. Software in the large. AT&T Technical Journal, 75(1):5–14, 1996. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bernstein:1997:SIS

[Ber97]  Lawrence Bernstein.  Soft-

ware investment strategy. Bell Labs Technical Journal, 2(3): 233–242, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Berkley:1990:HEH

[BF90] David A. Berkley and James L. Flanagan. Humanet: An experimental human-machine communications network based on ISDN wideband audio. AT&T Technical Journal, 69(5):87–99, September/October 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bielawski:1997:RPS

[BF97] William J. Bielawski III and David G. Furchtgott. 5ESS(R) packet switched network with ATM interconnect for CDMA. Bell Labs Technical Journal, 2 (2):203–212, Summer 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brunsen:1991:ASH

[BFF^{+}91] Lynn J. Brunsen, Arthur A. Frigo, Dennis J. Fitch, Frank K. Graff, and Thomas P. Groszczyk.

AT&T 5ESS switch hardware development methodology. a procedure for ensuring quality. AT&T Technical Journal, 70(2):63–72, 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Betker:1997:HM

[BFW97] Michael R. Betker, John S. Fernando, and Shaun P. Whalen. The history of the microprocessor. Bell Labs Technical Journal, 2(4):29–56, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bosco:1997:O

[BG97] Harry L. Bosco and Richard D.

Gitlin. Overview. Bell Labs Technical Journal, 2(2): 3–4, Summer 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Berg:1993:WCH

[BGH^{+}93] Mark R. Berg, Neil E. Gallensky, Greg D. Hamilton, Won U. Huh, James L. Knight, Jr., and John J. Sangermano. Worldclass hardware and transmission design. AT&T Technical Journal, 72(2):62–73, March/ April 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Borst:1997:WSS

[BGK^{+}97] Simon C. Borst, Sudheer A.

Grandhi, Colin L. Kahn, Krishnan Kumaran, Boris D.

Lubachevsky, and Donna M. Sand. Wireless simulation and self-organizing spectrum management. Bell Labs Technical Journal, 2(3):81–98, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bullard:1995:PDL

[BGM^{+}95] Glenn S. Bullard, Stanley Golinski, Brenda D. Morrison, William T. Myrick, Robert W. Neal, Jr., Roger L. Rose, Paul J. Rutkowski, Edward T. Schafer, and Walter H. Thurmond. 1B processor deployment: leading the way to flawless execution. AT&T Technical Journal, 74(3):61–72, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Blonder:1990:IPM

[BGT90] G. E. Blonder, R. A. Gottscho, and K. L. Tai. Interconnection processes and materials. AT&T Technical Journal, 69(6):46–59, November/December 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Benner:1991:AAS

[BH91] Robert E. Benner and Joe M. Harris. Applications, algorithms, and software for massively parallel computing. AT&T Technical Journal, 70 (6):59–72, November/December 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Benning:1997:ACM

[BHZ97] Roger D. Benning, M. Garry Hodgins, and George G. Zipfel Jr. Active control of mechanical vibrations. Bell Labs Technical Journal, 2(2):246–257, Summer 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bess:1992:DDP

[BJL92] C. Greg Bess, Walter W.

Jones, and Edward E. Lewis. Design and development of a program-and-place surfacemount-assembly machine. AT&T Technical Journal, 71(4):51–60, July/August 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Budka:1997:CDP

[BJS97] Kenneth C. Budka, Hong Jiang, and Steven E. Sommars. Cellular digital packet data networks. Bell Labs Technical Journal, 2 (3):164–181, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brooks:1994:ST

[BK94] Thomas A. Brooks and Michael M.

Kaplan. Security technologies. AT&T Technical Journal, 73(5): 4–8, September/October 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Buehrer:1999:IAW

[BKcL^{+}99] R. Michael Buehrer, Achilles G. Kogiantis, Shang chieh Liu, Jiann an Tsai, and Dirck Uptegrove. Intelligent antennas for wireless communications — uplink. Bell Labs Technical Journal, 4(3):73–103, Autumn 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Bohn:1992:FL

[BKNP92] Joseph M. Bohn, Peter P. Kania, Michael J. Nemchik, and Robert C. Purkey. Fiber in the loop. AT&T Technical Journal, 71(1):31–45, January/February 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Benson:1990:RLS

[BKP90] Kenneth E. Benson, Lionel C. Kimerling, and Peter T. Panousis. Reaching the limits in silicon processing. AT&T Technical Journal, 69(6):16–31, November/December 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bursh:1993:DRM

[BKR^{+}93] Talmage P. Bursh, Jr., Fredrick F. Kunzinger, LaJeana N. Roberts, Wesley L. Shanks, and Lisa A. Tantillo. Digital radio for mobile application. AT&T Technical Journal, 72(4):19–26, July/August 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bergland:1990:IFE

[BKSZ90] G. David Bergland, Geoffrey H.

Krader, D. Paul Smith, and Paul M. Zislis. Improving the front end of the softwaredevelopment process for largescale systems. AT&T Technical Journal, 69(2):7–21, 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Blaze:1994:IMT

[BLLR94] M. Blaze, J. Lacy, T. London, and M. Reiter. Issues and mechanisms for trustworthy systems: creating transparent mistrust. AT&T Technical Journal, 73(5):30–39, September/October 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Betta:1995:ESI

[BLR95] Carl E. Betta, John B. Lewis, and Larry A. Russell. The evolution of switch intelligence: an AT&T network perspective. AT&T Technical Journal, 74(3):10–17, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Becker:1994:SIN

[BLWW94] Daniel C. Becker, William J.

Liss, Jr., Arthur Weinstein, and Elizabeth A. Wilson. Star-Keeper II NMS — management of a cell relay network. AT&T Technical Journal, 73(4):46–55, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Brickell:1991:IID

[BM91] E. F. Brickell and K. S. McCurley. Interactive identification and digital signatures. AT&T Technical Journal, 70(6):73–86, November/December 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Blott:1999:ULB

[BMB^{+}99] Stephen M. Blott, Clifford E. Martin, Yuri J. Breitbart, José C. Brustoloni, Thomas R. Gramaglia, Henry F. Korth, David M. Kristol, Robert H. Liao, Eben L. Scanlon, and Avi Silberschatz. User-level billing and accounting in IP networks. Bell Labs Technical Journal, 4(4):237–251, October/December 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Benesty:1998:SSC

[BMHS98] Jacob Benesty, Dennis R. Morgan, Joseph L. Hall, and M. Mohan Sondhi. Synthesized stereo combined with acoustic echo cancellation for desktop conferencing. Bell Labs Technical Journal, 3(3):148–158, Autumn 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Berman:1994:CPS

[BMW94] Jay I. Berman, Harry H. Moore IV, and Jon R. Wright. Classic and prose stories: enabling technologies for knowledge-based systems. AT&T Technical Journal, 73(1):69–80, 1994. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Brinkman:1997:FSS

[BP97] William F. Brinkman and Mark R. Pinto. The future of solid-state electronics. Bell Labs Technical Journal, 2(4): 57–75, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Brownlie:1992:RTA

[BPP92] Robert Brownlie, James Prowse, and Madhav S. Phadke. Robust testing of AT&T PMX/ StarMAIL using OATS. AT&T Technical Journal, 71(3):41– 47, May/June 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Brown:1991:QEV

[Bro91] P. G. Brown. QFD: echoing the voice of the customer. AT&T Technical Journal, 70(2):18–32, 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Bohannon:1997:ADM

[BRSS97] Philip L. Bohannon, Rajeev R.

Rastogi, Avi Silberschatz, and S. Sudarshan. The architecture of the Dalí main memory storage manager. Bell Labs Technical Journal, 2(1):36–47, Spring 1997. CODEN BLTJFD.

ISSN 1089-7089 (print), 1538-7305 (electronic).

Bush:1990:CTM

[BRW90] M. Bush, K. Rasmussen, and Fai Wong. Conformance testing methodologies for OSI protocols. AT&T Technical Journal, 69(1):84–100, 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Buskens:1997:RMC

[BSP97] Richard W. Buskens, Muhammad A. Siddiqui, and Sanjoy Paul. Reliable multicasting of continuous data streams. Bell Labs Technical Journal, 2(2): 151–174, Summer 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Belanger:1992:RS

[BSW92] D. G. Belanger, E. E. Sumner, Jr., and P. J. Weinberger. Research in software. AT&T Technical Journal, 71(6):62–71, November/December 1992. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Busschbach:1998:TQC

[Bus98]  Peter B. Busschbach.  To-

ward QoS-capable virtual private networks. Bell Labs Technical Journal, 3(4):161–175, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Balakrishnan:1998:QDS

[BV98] Meera Balakrishnan and Ramanathan Venkateswaran. QoS and differentiated services in a multiservice network environment. Bell Labs Technical Journal, 3(4):222–239, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cheng:1991:MST

[CA91] Kwang-Ting Cheng and Vishwani D. Agrawal. Methods for synthesizing testable sequential circuits. AT&T Technical Journal, 70(1):64–86, January/February 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Crochiere:1995:DSP

[CB95] Ronald E. Crochiere and James R. Boddie. Digital signal processors: Toward Vision 2001. AT&T Technical Journal, 74(2):71–80, ???? 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cunningham:1993:GIN

[CBA93] J. P. Cunningham, C. D.

Blewett, and J. S. Anderson. Graphical interfaces for network operations and management. AT&T Technical Journal, 72(3): 57–66, May/June 1993. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Carlson:1996:EDO

[CBH96] Andrew C. Carlson, William R. Brook, and Christopher L. F. Haynes. Experiences with distributed objects. AT&T Technical Journal, 75(2):58–67, 1996. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Chen:1996:ITB

[CC96] Stephen Chen and Tsun S. Chow. Information technology — best practices in design and development. AT&T Technical Journal, 75(1):3–4, 1996. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Clougherty:1999:ARA

[CCJ^{+}99] Mark M. Clougherty, Donald E. Crowe, Kimerie W. Javitt, Fred C. Kemmerer, John Tardy, and John D. Unruh. The Any-Media(R) access system — providing solutions for distribution and network independence. Bell Labs Technical Journal, 4(2): 98–127, Summer 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Coffield:1995:LEH

[CCS95] Don R. Coffield, Robert G. Cole, and Vikram R. Saksena. The leading edge in highspeed data communications services: the Interspan Asynchronous Transfer Mode Service. AT&T Technical Journal, 74(4):38–49, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Crane:1994:TDD

[CDK^{+}94] F. Edward Crane III, Scott Davidson, Jack Kane, Charles E. Stroud, and Shianling Wu. Trends in digital device test methodologies. AT&T Technical Journal, 73(2):10–18, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Chen:1998:VMN

[CDT98] Zheng Chen, Kamala P. Das, and Benjamin Y. C. Tang. Voice and multiservice network design over ATM and IP networks. Bell Labs Technical Journal, 3(4):255–272, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Carr:1999:HIS

[CFH^{+}99] David C. Carr, Karen A. Folliard, Steel T. Huang, Ann M. Kenney, and Julie A. MacGregor. How to implement a successful communication program: A case study. Bell Labs Technical Journal, 4(3):172–180, Autumn 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Carboy:1990:SME

[CFJ^{+}90] James D. Carboy, George Foo, Lynn P. Jones, Lawrence E. Kinney, and Dan C. Krupka.

Striving for manufacturing excellence at the Denver Works: A summary. AT&T Technical Journal, 69(4):5–18, July/ August 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cleaveland:1996:DSP

[CFN96] J. Craig Cleaveland, Janet A. Fertig, and George W. Newsome. Dividing the software pie. AT&T Technical Journal, 75(2): 8–19, 1996. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Chen:1999:MON

[CFR^{+}99] Yi Chen, Mohammad T. Fatehi, Humberto J. La Roche, Jacob Z. Larsen, and Bruce L. Nelson. Metro optical networking. Bell Labs Technical Journal, 4(1): 163–186, Spring 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Castillo:1992:OOD

[CFW92] Cecilia M. Castillo, Elizabeth B. Flanagan, and Nancy W. Wilkinson. Object-oriented design and programming. AT&T Technical Journal, 71(6):27–33, November/December 1992. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cho:1990:PMP

[CG90] Alfred Y. Cho and Alastair M. Glass. Photonic materials and processing. AT&T Technical Journal, 69(6):77–91, November/December 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Carver:1994:NDO

[CGL^{+}94] G. E. Carver, M. L. Gray, J. Levkoff, B. W. Miller, and S. B. Phatak. Non-destructive optical techniques for characterizing semiconductor materials and devices. AT&T Technical Journal, 73(2):66–76, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Christenson:1996:EFC

[CH96] Dennis A. Christenson and Steel T. Huang. Estimating the fault content of software using the fix-on-fix model. Bell Labs Technical Journal, 1(1): 130–137, Spring 1996. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chen:1998:SCP

[Che98]  Xiaoqiang Chen.  Scalable

call processing architectures for ATM switches. Bell Labs Technical Journal, 3(1):62–80, Spring 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Crouch:1993:IPV

[CHJ93] Paul E. Crouch, J. Al Hicks, and John J. Jetzt. ISDN personal video. AT&T Technical Journal, 72(1):33–40, January/February 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Choy:1999:RMF

[CHKP99] Kwok-Chien Choy, Martin S. Halkyard, Shankar Krishnamoorthy, and Manas R. Panda. Reusable management frameworks for third-generation wireless networks. Bell Labs Technical Journal, 4(4):171–189, October/December 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chraplyvy:1999:HCL

[Chr99] Andrew R. Chraplyvy. Highcapacity lightwave transmission experiments. Bell Labs Technical Journal, 4(1):230–245, Spring 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chuah:1999:MVP

[CHV99] Mooi Choo Chuah and Enrique J. Hernandez-Valencia. Mobile virtual private dial-up services. Bell Labs Technical Journal, 4(3):51–72, Autumn 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chang:1999:MCS

[CJST99] Edward S. Chang, Aridaman K. Jain, David M. Slade, and S. Lee Tsao. Managing cyber security vulnerabilities in large networks. Bell Labs Technical Journal, 4(4):252–272, October/December 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Crabill:1994:SPS

[CK94] Edwin F. Crabill and James A.

Kukla. Service processing systems for AT&T’s intelligent network. AT&T Technical Journal, 73(6):39–47, 1994. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cox:1995:SCI

[CKC^{+}95] Richard V. Cox, Peter Kroon, Juin-Hwey Chen, Reed Thorkildsen, Kevin M. O’Dell, and David S. Isenberg. Speech coders: from idea to product. AT&T Technical Journal, 74(2):14–22, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cohen:1992:LTU

[CKL92] Melvin I. Cohen, Herwig Kogelnik, and Tingye Li. Lightwave technology: An update. AT&T Technical Journal, 71(1): 2–4, January/February 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Chambers:1998:VA

[CKLM98] Mahdi S. Chambers, Harbans Kaur, Terry G. Lyons, and Brian P. Murphy. Voice over ATM. Bell Labs Technical Journal, 3(4):176–191, October/December 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Clemens:1997:SMT

[Cle97] James T. Clemens. Silicon microelectronics technology. Bell Labs Technical Journal, 2(4): 76–102, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Culver-Lozo:1992:EAM

[CLG92] Kathleen Culver-Lozo and Vicki Glezman. Experiences applying methods supported by CASE tools. AT&T Technical Journal, 71(6):20–26, November/December 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Carter:1999:DRP

[CLN99] Galen L. Carter, Charles R. Lindemulder, and John M. Nichols. DEFINITY(R) ProLogix TM solutions rapid development. Bell Labs Technical Journal, 4 (3):181–190, Autumn 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cosky:1995:TMT

[CLRW95] Michael J. Cosky, Barry L. Lively, Linda A. Roberts, and Blake L. Wattenbarger. Talking to machines today and tomorrow: Designing for the user. AT&T Technical Journal, 74(2): 81–91, ???? 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Comizzoli:1990:RMP

[CLS90] Robert B. Comizzoli, James M.

Landwehr, and J. Douglas Sinclair. Robust materials and processes: Key to reliability. AT&T Technical Journal, 69(6):113– 128, November/December 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Choudhury:1995:EPM

[CLW95] Gagan L. Choudhury, Kin K.

Leung, and Ward Whitt. Efficiently providing multiple grades of service with protection against overloads in shared resources. AT&T Technical Journal, 74(4):50–63, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Chang:1993:WST

[CMS93] J. J. C. Chang, R. A. Miska, and R. A. Shober. Wireless systems and technologies: an overview. AT&T Technical Journal, 72(4): 11–18, July/August 1993. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Coplien:1991:IIS

[Cop91]  James O. Coplien.  ISH-

MAEL. an integrated software/ hardware maintenance and evolution environment. AT&T Technical Journal, 70(1):52–63, January/February 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Colson:1992:TQM

[CP92] J. S. Colson, Jr. and E. M. Prell. Total quality management for a large software project. AT&T Technical Journal, 71(3):48– 56, May/June 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Croak:1994:ETM

[Cro94] M. R. Croak. Emerging technologies and the messaging market. AT&T Technical Journal, 73(3):4–6, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Cohen:1990:SME

[CS90] Melvin I. Cohen and Laurence C. Seifert. Striving for manufacturing excellence. AT&T Technical Journal, 69(4): 2–4, July/August 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Carpenter:1996:NWB

[CS96] Jerry R. Carpenter and Michael J. Stima. New wireless business communications directions. Bell Labs Technical Journal, 1(2): 165–171, Summer 1996. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Cyr:1998:WBS

[CS98] Bernard L. Cyr and Michael R. Simmons. Wireless in-building services and architectures. Bell Labs Technical Journal, 3(1):

30–38, Spring 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chen:1997:EAI

[CT97a] Xiaoqiang Chen and Duke Thanes Tantiprasut. Evolution of ATM internetworking. Bell Labs Technical Journal, 2(2): 82–110, Summer 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chen:1997:VTM

[CT97b] Xiaoqiang Chen and Duke Thanes Tantiprasut. A virtual topology management protocol for ATM LAN emulation. Bell Labs Technical Journal, 2(2): 132–150, Summer 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Chih-Lin:1996:EMS

[CWH^{+}96] Chih-Lin I, Charles A. Webb III, Howard C. Huang, Stephan ten Brink, Sanjiv Nanda, and Richard D. Gitlin. IS-95 enhancements for multimedia services. Bell Labs Technical Journal, 1(2):60–87, Summer 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Deutschen:1996:AIS

[DBL96] Nelson R. Deutschen, E. Jay Bowers, and Jerry W. Lankford. ASCC: the impact of a silver bullet. AT&T Technical Journal, 75(1):24–34, 1996. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

DAnjou:1995:DPC

[DCGS95] Louis O. D’Anjou, Jae H. Choi, Werner J. Glantschnig, and Emil F. Stefanacci. Designing with plastics: considering part recyclability and the use of recycled materials. AT&T Technical Journal, 74(6):54–60, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Doshi:1999:OND

[DDH^{+}99] Bharat T. Doshi, Subrahmanyam Dravida, P. Harshavardhana, Oded Hauser, and Yufei Wang. Optical network design and restoration. Bell Labs Technical Journal, 4(1): 58–84, Spring 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Doshi:1998:CNG

[DDHQ98] Bharat T. Doshi, Subrahmanyam Dravida, P. Harshavardhana, and M. Akber Qureshi. A comparison of nextgeneration IP-centric transport architectures. Bell Labs Technical Journal, 3(4):63–85, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Doshi:1999:SDL

[DDHV^{+}99] Bharat T. Doshi, Subrahmanyam Dravida, Enrique J. Hernandez-Valencia, Wassim A. Matragi, M. Akber Qureshi, Jon Anderson, and James S. Manchester. A simple data link protocol for high-speed packet networks. Bell Labs Technical Journal, 4(1):85–104, Spring 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Doshi:1999:PAS

[DDJ^{+}99] Bharat T. Doshi, Subrahmanyam Dravida, Daniel R. Jeske, Kamala P. Murti, and Behrokh Samadi. Performance analysis of switch access systems. Bell Labs Technical Journal, 4(2):197–208, Summer 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Doshi:1996:BMA

[DDM^{+}96] Bharat T. Doshi, Subrahmanyam Dravida, Peter D. Magill, Curtis A. Siller Jr., and Kotikalapudi Sriram. A broadband multiple access protocol for STM, ATM, and variable length data services on hybrid fiber-coax networks. Bell Labs Technical Journal, 1(1):36–65, Spring 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dickinson:1995:GPM

[DDS^{+}95] David A. Dickinson, Clifton W.

Draper, Manjini Saminathan, John E. Sohn, and George Williams. Green product manufacturing. AT&T Technical Journal, 74(6):26–35, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Demetrescu:1999:LMA

[Dem99] Cristian Demetrescu. LLC-MAC analysis of general packet radio service in GSM. Bell Labs Technical Journal, 4(3): 37–50, Autumn 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dunlop:1997:MCI

[DER97] Alfred E. Dunlop, William J. Evans, and Lawrence A. Rigge. Managing complexity in IC design — past, present, and future. Bell Labs Technical Journal, 2(4):103–125, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

DAbadie:1992:FDP

[DFLN92] Catherine A. D’Abadie, Michael W. Funk, Francine S. Litofsky, and David E. Nelsen. The fast decision process. AT&T Technical Journal, 71(3):15–21, May/ June 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Denison:1999:NRN

[DG99] Brett A. Denison and Federico Giovannetti. NetCare(R) Network Management Services — managing multi-vendor networks. Bell Labs Technical Journal, 4(4):225–236, October/December 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dowden:1998:NGN

[DGM98] Douglas C. Dowden, Richard D.

Gitlin, and Robert L. Martin. Next-generation networks. Bell Labs Technical Journal, 3(4): 3–14, October/December 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Doshi:1998:PPC

[DHVS^{+}98] Bharat T. Doshi, Enrique J.

Hernandez-Valencia, Kotikalapudi Sriram, Yung-Terng Wang, and On-Ching Yue. Protocols, performance, and controls for voice over wide area packet networks. Bell Labs Technical Journal, 3(4):297–337, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

DeDuck:1992:FOL

[DJ92] Peter F. DeDuck and Steven R.

Johnson. The FT-2000 OC-48 Lightwave System. AT&T Technical Journal, 71(1):14–22, January/February 1992. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Doshi:1990:IPE

[DK90] Bharat T. Doshi and Dan C.

Krupka. Integration of planning and execution operations. theory and concepts. AT&T Technical Journal, 69(4):90–98, July/August 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Demaio:1993:BAS

[DK93a] M. Demaio and H. J. Kafka. Broadband applications and services prospectus. AT&T Technical Journal, 72(6):23–28, 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Dorato:1993:IDG

[DK93b] Karen H. Dorato and Lisa J. Klau. ISDN design for global markets. AT&T Technical Journal, 72(2):50–61, March/April 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Drakopoulos:1995:EEI

[DM95] E. Drakopoulos and M. J. Merges. Engineering and economic issues in managing the paradigm shift in computing environments. AT&T Technical Journal, 74(4):74–84, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Dowden:1995:IBL

[DMM^{+}95] Douglas C. Dowden, Robert E. McLear III, Alan H. Miller, Susan D. Schlough, David A. Welch, Ralph A. Wilson III, and Stephen A. Zeile. Improving on the best: “like a 1A, only better”. AT&T Technical Journal, 74(3):28–39, ???? 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

DAngelo:1994:SEM

[DMW94] Diana M. D’Angelo, Bruce Mc-Nair, and Joseph E. Wilkes. Security in electronic messaging systems. AT&T Technical Journal, 73(3):7–13, 1994. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Dorward:1997:ITO

[DPP^{+}97] Sean M. Dorward, Rob Pike, David Leo Presotto, Dennis M. Ritchie, Howard W. Trickey, and Philip Winterbottom. The Inferno^{T M} operating system. Bell Labs Technical Journal, 2 (1):5–18, Spring 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Dahbura:1990:AGP

[DS90] A. T. Dahbura and K. K. Sabnani. Algorithmic generation of protocol conformance tests. AT&T Technical Journal, 69 (1):101–118, 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ensor:1997:CMM

[EA97] J. Robert Ensor and Sudhir R.

Ahuja. Communication middleware for multi-party multimedia applications. Bell Labs Technical Journal, 2(1):74–85, Spring 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Eickmeyer:1995:NGU

[EET95] John R. Eickmeyer, Bryson E.

Epting, and Matthew V.

Thompson. The next generation in underwater acoustic detection. AT&T Technical Journal, 74(4):4–15, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Eick:1996:VCD

[EF96] S. G. Eick and D. E. Fyock. Visualizing corporate data. AT&T Technical Journal, 75(1):74–85, 1996. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Elwalid:1997:OMC

[EFG^{+}97] Anwar I. Elwalid, Glen G. Freundlich, Paul M. Gerhardt, Hassan Hagirahim, K. G. Ramakrishnan, and David Tse. An overview of the Multi-Media Communications eXchange (MMCX) and its performance characterization. Bell Labs Technical Journal, 2(2): 15–35, Summer 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Early:1993:VVT

[EKD93] S. H. Early, A. Kuzma, and E. Dorsey. The VideoPhone 2500 — video telephony on the public switched telephone network. AT&T Technical Journal, 72(1):22–32, January/February 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ejzak:1997:BSH

[EKK^{+}97] Richard P. Ejzak, Douglas N. Knisely, Sarath Kumar, Subhasis Laha, and Sanjiv Nanda. BALI: A solution for high-speed CDMA data. Bell Labs Technical Journal, 2(3):134–151, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ebner:1992:ATP

[ELC92] G. C. Ebner, T. K. Lybarger, and P. Coville. AT&T and telesystemes partnering in France. AT&T Technical Journal, 71(5):45–56, September/October 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Elliott:1992:PHA

[Ell92] W. Andrew Elliott. Progressive hand assembly of circuit packs in a small-lot, high-mix environment. AT&T Technical Journal, 71(4):45–50, July/August 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Elmendorf:1992:MQQ

[Elm92] David C. Elmendorf. Managing quality and quality improvement. AT&T Technical Journal, 71(3):57–65, May/June 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Erfani:1999:NME

[ELMS99] Shervin Erfani, Victor B.

Lawrence, Manu Malek, and Binay Sugla. Network management: Emerging trends and challenges. Bell Labs Technical Journal, 4(4):3–22, October/December 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Eicker:1991:IST

[EMS91] P. J. Eicker, D. J. Miller, and D. R. Strip. Intelligent systems and technologies for manufacturing. AT&T Technical Journal, 70(6):10–22, November/December 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Eng:1997:ASM

[EP97] Kai Y. Eng and Mark A. Pashan. Advances in sharedmemory designs for gigabit ATM switching. Bell Labs Technical Journal, 2(2):175– 187, Summer 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Evans:1996:LCR

[ESWW96] James G. Evans, R. Anthony Shober, Stephen A. Wilkus, and Gregory A. Wright. A low-cost radio for an electronic price label system. Bell Labs Technical Journal, 1(2):203–215, Summer 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Farley:1992:SUM

[Far92] D. L. Farley. The 5ESS switch US market front end process. AT&T Technical Journal, 71(3): 8–14, May/June 1992. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Farber:1993:UTA

[Far93] J. M. Farber. Using technology to achieve ease of use. AT&T Technical Journal, 72 (3):4–6, May/June 1993. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Furman:1999:SBS

[FCT^{+}99] Daniel S. Furman, Michael J.

Cosky, David L. Thomson, Stephen A. O’Brien, and Eric E. Sumner Jr. Speech-based services. Bell Labs Technical Journal, 4(2):88–97, Summer 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Flanagan:1990:SPP

[FD90] J. L. Flanagan and C. J. Del Riesgo. Speech processing: a perspective on the science and its applications. AT&T Technical Journal, 69(5):2–13, September/October 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Feiner:1991:DPA

[FE91] Alec Feiner and Katherine A.

Edward. Development processes and applications at AT&T. an overview. AT&T Technical Journal, 70(2):2–6, 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Feiner:1992:OQP

[FE92] Alec Feiner and Ketherine A. Edward. An overview of quality and project management processes at AT&T. AT&T Technical Journal, 71(3):2–7, May/ June 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Foo:1992:VCF

[FF92] George Foo and David J. Friedman. Variability and capability: The foundation of competitive operations performance. AT&T Technical Journal, 71(4): 2–9, July/August 1992. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fenton:1993:DDG

[FFM93] Frank M. Fenton, Kenneth T. Fong, and Albert Mizrahi. Design and development for a global market. AT&T Technical Journal, 72(2):2–5, March/ April 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Faynberg:1997:DWI

[FGJL97] Igor Faynberg, Lawrence R. Gabuzda, Terry Jacobson, and Hui-Lan Lu. The development of the wireless intelligent network (WIN) and its relation to the international intelligent network standards. Bell Labs Technical Journal, 2(3):57–80, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fallah:1991:DPA

[FHM^{+}91] M. H. Fallah, J. P. Holtman, J. F. Maranzano, D. P. Smith, and G. T. Tucker. Development process audits and reviews. AT&T Technical Journal, 70(2):99–108, 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Flora-Holmquist:1997:VFS

[FHMOS97] Alan R. Flora-Holmquist, Edward Morton, James D. O’Grady, and Mark G. Staskauskas. The virtual finite-state machine design and implementation paradigm. Bell Labs Technical Journal, 2(1):96–113, Spring 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fowler:1992:TTB

[FHO92] Glenn S. Fowler, James E.

Humelsine, and Carl H. Olson. Tools and techniques for building and testing software systems. AT&T Technical Journal, 71(6):46–61, November/December 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fallah:1994:SPA

[FJ94] M. Hosein Fallah and Ahmad M. Jrad. SQA — a proactive approach to assuring software quality. AT&T Technical Journal, 73(1):26–33, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fry:1996:NGW

[FJL^{+}96] George E. Fry, Albert Jordan, David Y. Lee, Anil S. Sawkar, Nitin J. Shah, and William C. Wiberg. Next generation wireless networks. Bell Labs Technical Journal, 1(2):88–96, Summer 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fossett:1992:ERC

[FKCR92] L. D. Fossett, A. Kasper, S. Civaniar, and W. D. Radwill. Evolution of routing and capacity management in international switched networks. AT&T Technical Journal, 71 (5):13–23, September/October 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fischell:1990:IVT

[FKF90] David R. Fischell, Sarbmeet S. Kanwal, and Daniel Furman. Interactive voice technology applications. AT&T Technical Journal, 69(5):61–76, September/October 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Frisch:1993:ICP

[FKK93] Howard D. Frisch, Bryan S. Katz, and James L. Knight. International considerations in PBX design. AT&T Technical Journal, 72(2):17–23, March/ April 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fendick:1998:PTI

[FKLS98] Kerry W. Fendick, Vijay P.

Kumar, T. V. Lakshman, and Dimitrios Stiliadis. The PacketStar^{T M}6400 IP switch — an IP switch for the converged network. Bell Labs Technical Journal, 3(4):32–47, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Farrell:1992:ISE

[FKRW92] Brian L. Farrell, William Kahan, Jay W. Reightler, and Alan C. Woodard. Improving service efficiency in manufacturing integrated circuits. AT&T Technical Journal, 71(4):37–44, July/August 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Funka-Lea:1998:QRP

[FLJL^{+}98] Cynthia A. Funka-Lea, Carol L.

Janczewski, Wing Cheong Lau, Ramesh Nagarajan, Yung-Terng Wang, and Zhi-Lei Xin. QoS routing and performance in packet networks: A visual simulation platform and case study. Bell Labs Technical Journal, 3(4):240–254, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

FitzGerald:1995:AIR

[FM95] Cary W. FitzGerald and John P. Moosmiller. Architecture of the Intuity Response Application Programming Interface (IRAPI). AT&T Technical Journal, 74(2):92–101, ???? 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Flynn:1995:MIP

[FO95] Lisa A. Flynn and Thomas M. O’Leary. Manufacturing issues of the 1B processor. AT&T Technical Journal, 74(3):49–60, ???? 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Foster:1991:CCA

[Fos91] Jeffrey C. Foster. CAE/CAD at AT&T: An introduction. AT&T Technical Journal, 70(1):2–8, January/February 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Foschini:1996:LST

[Fos96] Gerard J. Foschini. Layered space-time architecture for wireless communication in a fading environment when using multi-element antennas. Bell Labs Technical Journal, 1(2): 41–59, Summer 1996. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Fenton:1996:ATT

[FS96] Frank M. Fenton and James D. Sipes. Architectural and technological trends in access: An overview. Bell Labs Technical Journal, 1(1):3–10, Spring 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Foard:1991:SCN

[FSV91] Christopher F. Foard, Leslie A.

Spindel, and Pramode K. Verma. Switch-to-computer networking in the nineties: The evolution of AT&T’s switchcomputer interface. AT&T Technical Journal, 70(5):45–58, September/October 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fossaceca:1998:PTA

[FSW98] John M. Fossaceca, James D.

Sandoz, and Philip Winterbottom. The PathStar^{T M} access server: Facilitating carrier-scale packet telephony. Bell Labs Technical Journal, 3(4):86–102, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Garg:1999:TMC

[Gar99] Vijay K. Garg. Trunk management in cellular/PCS networks. Bell Labs Technical Journal, 4 (3):120–133, Autumn 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gould:1993:RSC

[GCJM93] Kevin W. Gould, Richard V.

Cox, Nikil S. Jayant, and Melvin J. Melchner. Robust speech coding for the indoor wireless channel. AT&T Technical Journal, 72(4):64–73, July/August 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Graedel:1995:GPD

[GCS95] Thomas E. Graedel, Paul Reaves Comrie, and Janine C. Sekutowski. Green product design. AT&T Technical Journal, 74(6):17–25, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gelman:1994:SBI

[Gel94] Stacey Gelman. Silver bullet: an iterative model for process definition and improvement. AT&T Technical Journal, 73(4):35–45, 1994. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gatenby:1990:DXD

[GF90] D. A. Gatenby and G. Foo. Design for X (DFX): key to competitive, profitable products. AT&T Technical Journal, 69(3): 2–13, 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Graedel:1990:IEI

[GG90] Thomas E. Graedel and Leslie A. Guth. The impact of environmental issues on materials and processes. AT&T Technical Journal, 69(6):129–140, November/December 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gana:1997:SMA

[GH97] Akli Gana and Steel T. Huang.

Statistical modeling applied to managing global 5ESS(R)-2000 switch software development. Bell Labs Technical Journal, 2 (1):144–153, Spring 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Godefroid:1998:SST

[GHJ98] Patrice Godefroid, Robert S.

Hanmer, and Lalita Jategaonkar Jagadeesan. Systematic software testing using VeriSoft — An analysis of the 4ESS^{T M} heart-beat monitor. Bell Labs Technical Journal, 3 (2):32–46, Summer 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Geusic:1992:QWL

[GHK^{+}92] J. E. Geusic, R. L. Hartman, Uziel Koren, Won-Tien Tsang, and D. P. Wilt. Quantum well lasers in telecommunications. AT&T Technical Journal, 71(1):75–83, January/February 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gauldin:1993:WMS

[GHK^{+}93] Mark A. Gauldin, Jon H. Horwedel, Jack P. Kozik, Robin J. Thompson, and Randall J. Wilson. 5ESS wireless mobile switching center. AT&T Technical Journal, 72(4):38–47, July/August 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Goudreault:1999:IIA

[GHN^{+}99] Robert A. Goudreault, P. Harshavardhana, Ramesh Nagarajan, Carol Y. Sunada-Wong, and Mark R. Wilson. Impact of an integrated architecture for bandwidth management in a broadband network infrastructure. Bell Labs Technical Journal, 4(1):19–41, Spring 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gregus:1998:CSM

[GLDT98] Jeffrey A. Gregus, Maureen Y. Lau, Yinon Degani, and King L. Tai. Chip-scale modules for high-level integration in the 21 st century. Bell Labs Technical Journal, 3(3):116–124, Autumn 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gatenby:1994:CEE

[GLH^{+}94] David A. Gatenby, Paul M. Lee, Randall E. Howard, Kaveh Hushyar, Rich Layendecker, and John Wesner. Concurrent engineering: an enabler for fast, high-quality product realization. AT&T Technical Journal, 73(1):34–47, 1994. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gelman:1992:CLS

[GLM92] Stacey J. Gelman, Fred M. Lax, and Joseph F. Maranzano. Competing in large-scale software development. AT&T Technical Journal, 71(6):2–11, November/December 1992. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Grzelak:1996:RDS

[GMSW96] Stephen A. Grzelak, Harrison Miles Jr., Edward S. Szurkowski, and William P. Weber Jr. Residential data services via hybrid fiber-coax local access networks. Bell Labs Technical Journal, 1(1):88–99, Spring 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gagen:1996:HFC

[GP96a] Paul F. Gagen and William E.

Pugh. Hybrid fiber-coax access networks. Bell Labs Technical Journal, 1(1):28–35, Spring 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gelman:1996:BBI

[GP96b] Stacey J. Gelman and W. Douglas Peck. Bringing business information to AT&T network systems through a data warehouse. AT&T Technical Journal, 75(2):68–78, 1996. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Grimes:1991:SID

[GRL^{+}91] Gary J. Grimes, Mark S.

Russo, Christopher Lanzafame, Christopher D. Near, and Bryan S. Moffitt. Synchronization in intelligent digital networks. AT&T Technical Journal, 70(5):59–68, September/October 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Glass:1992:MOH

[GS92] K. K. Glass and L. M. Sanders. Management organizational handoffs with empowered teams. AT&T Technical Journal, 71(3):22–30, May/ June 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Garg:1997:TNA

[GS97] Vijay K. Garg and Elbert L. Sneed. TDMA for North American PCS systems. Bell Labs Technical Journal, 2(3): 207–220, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Giles:1999:WAD

[GS99] C. Randy Giles and Magaly Spector. The wavelength add/ drop multiplexer for lightwave communication networks. Bell Labs Technical Journal, 4(1): 207–229, Spring 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Goyal:1998:DBC

[GUP98] Suresh Goyal, Sanjay S. Upasani, and Dhiren M. Patel. Designing best-in-class impacttolerant cellular phones and other portable products. Bell Labs Technical Journal, 3(3): 159–174, Autumn 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Greenspan:1992:PDT

[GWB92] Steven L. Greenspan, John J.

Wisowaty, and Raymond E. Bright, Jr. Product definition through the design-by-use process. AT&T Technical Journal, 71(3):31–40, May/June 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Gitlin:1996:EWW

[GZ96] Richard D. Gitlin and George I.

Zysman. The expanding world of wireless technology. Bell Labs Technical Journal, 1(2): 3–6, Summer 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Gitlin:1997:O

[GZ97] Richard D. Gitlin and George I.

Zysman. Overview. Bell Labs Technical Journal, 2(3): 3–4, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hornbach:1993:SNG

[HBCS93] Barbara H. Hornbach, William J.

Bielawski, Ronald J. Canniff, and Paul A. Stiling. 5ESS-2000 switch: The next generation switching system. AT&T Technical Journal, 72(5):4–13, September/October 1993. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Harrison:1996:PPS

[HC96] Neil B. Harrison and James O. Coplien. Patterns of productive software organizations. Bell Labs Technical Journal, 1(1): 138–145, Spring 1996. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Heirman:1993:IMP

[HCS93] Donald N. Heirman, Steven M. Crosby, and Douglas C. Smith. International mandatory product standards. AT&T Technical Journal, 72(2):25–40, March/ April 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hall:1991:ASC

[HCWW91] Harry M. Hall, Christopher D. Carson, Earle H. West, and Lonnie D. Whitehead. The AT&T service ciruit node: A new element for providing intelligent network services. AT&T Technical Journal, 70(3–4):72– 84, March/April 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Horenkamp:1993:SCS

[HE93] John J. Horenkamp and J. Scott Eudy. Structured cabling systems for the global premises distribution market. AT&T Technical Journal, 72(2):41–49, March/April 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

He:1998:CIN

[HE98] Taosong He and Stephen G.

Eick. Constructing interactive network visual interfaces. Bell Labs Technical Journal, 3(2): 47–57, Summer 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hendrix:1990:MET

[Hen90] Stanley A. Hendrix. Manufacturing execution: Theory and concepts. AT&T Technical Journal, 69(4):33–52, July/August 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Holland:1994:SEC

[HHJ^{+}94] Hans G. Holland, C. P. Huang, Charles E. Jeschke, Robert A. Pergande, Lee C. Stecher, and David J. Stelte. The 5ESS-2000 switch: exceeding customer expectations. AT&T Technical Journal, 73(6):28–38, 1994. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hsueh:1994:SPA

[HHMP94] T. Richard Hsueh, Thomas F.

Houghton, Joseph F. Maranzano, and Gerald P. Pasternack. Software production: from art/ craft to engineering. AT&T Technical Journal, 73(1):59–68, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Harvey:1991:CCS

[HHP91] Dean E. Harvey, Shannon M. Hogan, and John Y. Payseur. Call center solutions. AT&T Technical Journal, 70(5):36–44, September/October 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hansen:1997:CES

[HJ97] Mark H. Hansen and David A. James. A computing environment for spatial data analysis in the microelectronics industry. Bell Labs Technical Journal, 2 (1):114–129, Spring 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hassler:1995:RDC

[HJKN95] Kerry W. Hassler, Cynthia C. Jones, Joylee E. Kohler, and Robert D. Nalbone. Revolutionizing DEFINITY call centers in the 1990s. AT&T Technical Journal, 74(4):64–73, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hermann:1995:ECA

[HK95] Ferdinand Hermann and Peter Kleinsorge. Environmental certification of ATT-GIS in Augsburg, Germany. AT&T Technical Journal, 74(6):72–78, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Huang:1996:CSF

[HKBW96] Yennun Huang, Chandra M. R.

Kintala, Lawrence Bernstein, and Yi-Min Wang. Components for software fault tolerance and rejuvenation. AT&T Technical Journal, 75(2):29–37, 1996. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Horn:1993:SBM

[HLMS93] D. N. Horn, T. G. Lyons, J. C. Mitchell, and D. L. Skran. A standards-based multimedia conferencing bridge. AT&T Technical Journal, 72(1):41–49, January/February 1993. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hinton:1992:FSP

[HM92] H. Scott Hinton and David A. B.

Miller. Free-space photonics in switching. AT&T Technical Journal, 71(1):84–92, January/February 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Heirman:1990:DME

[HMC90] D. N. Heirman, R. C. Morris, and S. M. Crosby. Design for multinational EMC compliance. AT&T Technical Journal, 69(3):28–45, 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hillis:1995:IAP

[HMD^{+}95] Margaret A. Hillis, Gail W. Minkus, Craig L. DeCaluwe, Michael N. Meyers, and Edward T. Schafer. An improved approach to product quality through testing. AT&T Technical Journal, 74(3):40–48, ???? 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ho:1999:RTP

[HMH99] L. Lawrence Ho, Christopher J. Macey, and Ronald G. Hiller. Real-time performance monitoring and anomaly detection in the Internet: An adaptive, objective-driven, mix-and-match approach. Bell Labs Technical Journal, 4(4):23–41, October/December 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hayden:1990:SDA

[HMMS90] C. C. Hayden, J. C. Mitchell, J. Mukerji, and F. A. Schmidt. The Software Development Assistant. AT&T Technical Journal, 69(2):76–89, 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Holzmann:1990:AAP

[Hol90] G. J. Holzmann. Algorithms for automated protocol verification. AT&T Technical Journal, 69(1):32–44, 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Holzmann:1997:DTR

[HPR97] Gerard J. Holzmann, Doron A.

Peled, and Margaret H. Redberg. Design tools for requirements engineering. Bell Labs Technical Journal, 2(1):86–95, Spring 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Henneman:1993:NDT

[HR93] Richard L. Henneman and David M. Rubini. New directions in transaction terminal interfaces. AT&T Technical Journal, 72(3):50–56, May/ June 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hirschberg:1990:VRS

[HRRS90] Julia B. Hirschberg, Stephen A.

Riederer, James E. Rowley, and Ann K. Syrdal. Voice response systems: Technologies and applications. AT&T Technical Journal, 69(5):42–51, September/October 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Hsu:1995:ASE

[HS95] Jiunn Carl Hsu and Larry A.

Seese. AT&T switching evolution challenge. AT&T Technical Journal, 74(3):6–9, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Holzmann:1998:IRT

[HS98] Gerard J. Holzmann and Margaret H. Smith. Interval reduction through requirements analysis. Bell Labs Technical Journal, 3(2):22–31, Summer 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hossain:1999:PBN

[HSGR99] Ashfaq Hossain, Houshing F. Shu, Charles R. Gasman, and Randolph A. Royer. Policybased network load management. Bell Labs Technical Journal, 4(4):95–108, October/December 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

He:1998:CAC

[HW98] Linhai He and Albert K. Wong. Connection admission control design for GlobeView(R)-2000 ATM core switches. Bell Labs Technical Journal, 3(1):94–111, Spring 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Hlavacek:1997:AMI

[HZ97] Donna M. Hlavacek and Robert J. Zurawski. Alternative methods for introducing new wireless intelligent network services using triggers and queries. Bell Labs Technical Journal, 2(3): 20–29, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Infosino:1991:MAP

[IPU91] William J. Infosino, Sam H. Parker, and David G. Unger. Market analysis and product design for telecommunications equipment and services. AT&T Technical Journal, 70(2):7–17, 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Joyce:1991:MDC

[JAC91] John A. Joyce, Marc C. Ayres, and Robert F. Cruickshank, III. Managing design changes. AT&T Technical Journal, 70 (3–4):73–83, March/April 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Jackman:1998:MOF

[Jac98] Neil A. Jackman. Multilayer optical filters with random errors. Bell Labs Technical Journal, 3 (1):112–121, Spring 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jayant:1995:MTD

[JALR95] Nikil Jayant, Bryan D. Ackland, Victor B. Lawrence, and Lawrence R. Rabiner. Multimedia: technology dimensions and challenges. AT&T Technical Journal, 74(5):14–33, September/October 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Jayant:1995:ACT

[JC95] Nikil S. Jayant and Edward Y.

Chen. Audio compression: technology and applications. AT&T Technical Journal, 74(2):23–34, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Jayant:1990:CSW

[JLP90] N. S. Jayant, V. B. Lawrence, and D. P. Prezas. Coding of speech and wideband audio. AT&T Technical Journal, 69 (5):25–41, September/October 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Johnson:1999:AON

[JN99] Steven R. Johnson and Virginia L. Nichols. Advanced optical networking — Lucent’s MONET network elements. Bell Labs Technical Journal, 4(1): 145–162, Spring 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Jackman:1999:OCC

[JPMK99] Neil A. Jackman, Sunita H. Patel, Benny P. Mikkelsen, and Steven K. Korotky. Optical cross connects for optical networking. Bell Labs Technical Journal, 4(1):262–281, Spring 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Juang:1995:DAS

[JPT95] Biing-Hwang Juang, Robert J. Perdue, Jr., and David L. Thomson. Deployable automatic speech recognition systems: advances and challenges. AT&T Technical Journal, 74 (2):45–56, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Jackel:1995:OCR

[JSS^{+}95] Lawrence D. Jackel, David Sharman, Charles E. Stenard, B. Ivan Strom, and Derek Zuckert. Optical character recognition for self-service banking. AT&T Technical Journal, 74(4):16–24, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Johnson:1996:USS

[JW96] Graham I. Johnson and Mark G.

Westwater. Usability and selfservice information technology: cognitive engineering in product design and evaluation. AT&T Technical Journal, 75(1):64–73, 1996. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Jorgensen:1991:RHM

[JY91] James L. Jorgensen and David S.

Yaney. Radiation-hardened microelectronics. AT&T Technical Journal, 70(6):31–36, November/December 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Khoe:1999:DLC

[KBB99] Hoo-Yin Khoe, A. Craig Bolling, and Arupjyoti Bhuyan. Digital loop carrier solutions for voice and data networks. Bell Labs Technical Journal, 4(2): 209–217, Summer 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kruisbrink:1997:SDE

[KBP97] J. Casper Kruisbrink, Jaap Bood, and Mary Ann Parsons. On-site data evolution for a 5ESS(R) switch retrofit. Bell Labs Technical Journal, 2 (1):65–73, Spring 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Khetarpal:1997:SSM

[KD97] Rakesh Khetarpal and S. Krishna Dronamraju. SMART — a solution for managing the capacity, growth, and performance of wireless networks. Bell Labs Technical Journal, 2(3): 182–193, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kinney:1990:HOM

[KDF90] Lawrence E. Kinney, Elizabeth T. Denning, and George Foo. High-option manufacturing: The Denver Works Information System Architecture. AT&T Technical Journal, 69 (4):110–116, July/August 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kordahi:1995:IMT

[KGC95] Maurice E. Kordahi, Robert F. Gleason, and Ta-Mu Chien. Installation and maintenance technology for undersea cable systems. AT&T Technical Journal, 74(1):60–74, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Khanna:1997:PTA

[Kha97] Sanjeev Khanna. A polynomial time approximation scheme for the SONET ring loading problem. Bell Labs Technical Journal, 2(2):36–41, Summer 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kowalski:1990:IEO

[KHD90] T. J. Kowalski, Yean-Ming Huang, and H. V. Diamantidis. An interpretive environment for operations support systems. AT&T Technical Journal, 69(2):42–51, 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kinney:1990:MP

[Kin90] Lawrence E. Kinney. Material planning. AT&T Technical Journal, 69(4):19–32, July/August 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kriaras:1997:TGM

[KJPR97] Ioannis N. Kriaras, André W.

Jarvis, Vincent E. Phillips, and Derek J. Richards. Thirdgeneration mobile network architectures for the Universal Mobile Telecommunications System (UMTS). Bell Labs Technical Journal, 2(3):99– 117, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Koch:1992:PIC

[KK92] Thomas L. Koch and Uziel Koren. Photonic integrated circuits. AT&T Technical Journal, 71(1):63–74, January/February 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kirova:1998:DCS

[KK98] Vassilka D. Kirova and Howard G. Kradjel. The DirSA case study: An introduction to software architecture technology. Bell Labs Technical Journal, 3(3): 125–138, Autumn 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Knisely:1998:CTG

[KLR98] Douglas N. Knisely, Quinn Li, and Nallepilli S. Ramesh. cdma2000: A third-generation radio transmission technology. Bell Labs Technical Journal, 3 (3):63–78, Autumn 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kamerman:1997:WRI

[KM97] Ad Kamerman and Leo Monteban. WaveLAN(R)-II: a highperformance wireless LAN for the unlicensed band. Bell Labs Technical Journal, 2(3): 118–133, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kozik:1998:VSN

[KMS98] Jack Kozik, Warren A. Montgomery, and John J. Stanaway Jr. Voice services in nextgeneration networks: The evolution of the intelligent network and its role in generating new revenue opportunities. Bell Labs Technical Journal, 3(4):124–143, October/December 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kettler:1992:AGI

[KNS^{+}92] H. W. Kettler, G. Natarajan, E. W. Scher, P. Y. Shih, and P. Marr Wainscott. AT&T’s global intelligent network architecture. AT&T Technical Journal, 71(5):30–35, September/October 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kerfoot:1995:FDU

[KR95] Franklin W. Kerfoot and Peter K. Runge. Future directions for undersea communications. AT&T Technical Journal, 74(1):93–102, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kowalik:1998:DBT

[KRS98] John Kowalik Jr., John J.

Rosinski, and Bradford R. Siepman. Digital business telephones — project laser. Bell Labs Technical Journal, 3(1): 122–133, Spring 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kocan:1998:UPA

[KRZ98] Kristin F. Kocan, Patricia A. Riley, and Robert J. Zurawski. Universal premises access to voice band and ATM networks using a broadband interworking gateway. Bell Labs Technical Journal, 3(2):93–108, Summer 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Kamath:1993:RBO

[KSS93] Yogeesh H. Kamath, Ruth E. Smilan, and Jean G. Smith. Reaping benefits with objectoriented technology. AT&T Technical Journal, 72(5):14–24, September/October 1993. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kythe:1996:PDB

[Kyt96] Dave K. Kythe. The promise of distributed business components. AT&T Technical Journal, 75(2):20–28, 1996. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Lee:1997:APA

[LB97] Anne Y. Lee and Bohdan L. Bodnar. Architecture and performance analysis of packetbased mobile switching centerto-base station traffic communications for TDMA. Bell Labs Technical Journal, 2(3):

46–56, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

LuValle:1998:SEA

[LCK^{+}98] Michael J. LuValle, Leon R.

Copeland, Sriraman Kannan, Justin B. Judkins, and Paul J. Lemaire. A strategy for extrapolation in accelerated testing. Bell Labs Technical Journal, 3 (3):139–147, Autumn 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Layadi:1999:IPE

[LCL99] Nace Layadi, Jennifer I. Colonell, and John Tseng-Chung Lee. An introduction to plasma etching for VLSI circuit technology. Bell Labs Technical Journal, 4(3): 155–171, Autumn 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lee:1993:SBL

[LGPR93] Chin-Hui Lee, Jean-Luc Gauvain, Roberto Pieraccini, and Lawrence R. Rabiner. Subwordbased large-vocabulary speech recognition. AT&T Technical Journal, 72(5):25–36, September/October 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Liss:1995:NPO

[LK95] Jonathan M. Liss and Kathleen A. Kurek. Network planning, operation, and maintenance practices for undersea communication systems. AT&T Technical Journal, 74(1):75–82, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Libman:1995:IVN

[LMMN95] Roger E. Libman, Mowaffak T. Midani, Irene J. Morgan, and Hung Tan Nguyen. The interactive video network: an overview of the Video Manager and the V Protocol. AT&T Technical Journal, 74(5):92–105, September/October 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Lanzerotti:1995:SLS

[LMMT95] Louis J. Lanzerotti, Lester V. Medford, Carol G. Maclennan, and David J. Thomson. Studies of large-scale Earth potentials across oceanic distances. AT&T Technical Journal, 74(3):73–84, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Laudise:1990:CRP

[LMN90] R. A. Laudise, D. W. McCall, and S. R. Nagel. The critical role of processing in leadingedge product capability. AT&T Technical Journal, 69(6):9–15, November/December 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Lie:1992:GTD

[LMN^{+}92] Gwan K. Lie, Thomas L. McRoberts, Jan Neuteboom, Robert E. Schleicher, and Jeffrey D. Wiggs. Global teamwork: Developing international ISDN capabilities for the 5ESS switch. AT&T Technical Journal, 71(5):24–29, September/October 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Lynch:1995:DDO

[LMW95] Robert L. Lynch, Richard L.

Maybach, and Philip F. Walch. Design and deployment of optically amplified undersea systems. AT&T Technical Journal, 74(1):83–92, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Lied:1997:ISR

[LPH97] Robert Lied, Lynn P. Pautler, and Patrick E. Helmers. Introducing software reuse technology. Bell Labs Technical Journal, 2(1):188–199, Spring 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lin:1998:NIM

[LPS98] Geng Lin, John D. Price, and Tanjore K. Srinivas. Network information models and OneVision(R) architecture. Bell Labs Technical Journal, 3(4):208– 221, October/December 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lakshmi-Ratan:1999:LTS

[LR99] Ramnath A. Lakshmi-Ratan.

The Lucent Technologies Softswitch — Realizing the promise of convergence. Bell Labs Technical Journal, 4(2):174–195, Summer 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Luss:1990:IPE

[LRW90] Hanan Luss, Moshe B. Rosenwein, and Elizabeth T. Wahls. Integration of planning and execution operations: Finalassembly sequencing. AT&T Technical Journal, 69(4):99– 109, July/August 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Lawrence:1996:BAH

[LSWZ96] Victor B. Lawrence, Luke J. Smithwick, Jean-Jacques Werner, and Nikolaos A. Zervos. Broadband access to the home on copper. Bell Labs Technical Journal, 1(1):100–114, Spring 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Lanzerotti:1997:WHA

[LTM97] Louis J. Lanzerotti, David J. Thomson, and Carol G. Maclennan. Wireless at high altitudes — environmental effects on space-based assets. Bell Labs Technical Journal, 2(3): 5–19, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Martin:1997:DNS

[MAB^{+}97] Samuel Martin, Vance D.

Archer III, David M. Boulin, Michel R. Frei, Kwok K. Ng, and Ran-Hong Yan. Device noise in silicon RF technologies. Bell Labs Technical Journal, 2 (3):30–45, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Martin:1997:O

[Mar97] Robert L. Martin. Overview.

Bell Labs Technical Journal, 2 (1):3–4, Spring 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mayer:1991:SNI

[MASY91] Robert L. Mayer, Ronald G.

Appleton, E. Lee Sneed, and Stanley Y. Yeh. Service Net-2000: An intelligent network evolution. AT&T Technical Journal, 70(3–4):99–110, March/April 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Maxemchuk:1994:EDD

[Max94]  N. F. Maxemchuk.  Elec-

tronic document distribution. AT&T Technical Journal, 73 (5):73–80, September/October 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic). URL http:

//www.cl.cam.ac.uk/~fapp2/
steganography/bibliography/
034157.html.

Mayo:1992:PSE

[May92] J. S. Mayo. Partnering: a strategic element of globalization [international telecommunications]. AT&T Technical Journal, 71(5):2–5, September/October 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Miller:1991:DNS

[MC91] John A. Miller and Chung Ping Cheng. Designing networking solutions for the nineties: A new approach. AT&T Technical Journal, 70(5):14–26, September/October 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Morgan:1991:SCT

[MCG^{+}91] Michael J. Morgan, Michael J. Cosky, Thomas M. Gruenenfelder, T. Curtis Holmes, Jr., and Gerald A. Raack. Service creation technologies for the intelligent network. AT&T Technical Journal, 70(3–4):58– 71, March/April 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

McGrath:1992:LT

[McG92] Carl J. McGrath. AM lightwave technology. AT&T Technical Journal, 71(1):23–30, January/February 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Miyoshi:1991:STB

[MG91] Dennis S. Miyoshi and Mary W.

Green. Security and technology. A better mousetrap. AT&T Technical Journal, 70(6):87–91, November/December 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Midha:1997:SCM

[Mid97] Anil K. Midha. Software configuration management for the 21st Century. Bell Labs Technical Journal, 2(1):154–165, Spring 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Millen:1993:PBU

[Mil93] David R. Millen. Pen-based user interfaces. AT&T Technical Journal, 72(3):21–27, May/ June 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Mitra:1994:IAM

[Mit94] Nilotpal Mitra. Introduction to the ASN.1 macro replacement notation. AT&T Technical Journal, 73(3):66–79, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Meyer:1996:WES

[MJP^{+}96] Michael J. Meyer, Terry Jacobson, Maria E. Palamara, Elizabeth A. Kidwell, Robert E. Richton, and Giovanni Vannucci. Wireless enhanced 9-1-1 service — making it a reality.

Bell Labs Technical Journal, 1 (2):188–202, Summer 1996. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mortenson:1995:UOA

[MJSS95] Robert L. Mortenson, B. Scott Jackson, Seymour Shapiro, and William F. Sirocky. Undersea optically amplified repeatered technology, products, and challenges. AT&T Technical Journal, 74(1):33–46, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Moghe:1998:TQA

[MK98] Pratyush Moghé and Asawaree Kalavade. Terminal QoS of adaptive applications. Bell Labs Technical Journal, 3(2): 76–92, Summer 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Manzione:1990:PEP

[ML90] Louis T. Manzione and David J. Lando. Polymers in electronics packaging. AT&T Technical Journal, 69(6):60–76, November/December 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

McKinney:1998:SCN

[MMO^{+}98] R. D. McKinney, Warren A. Montgomery, Hamza Ouibrahim, Paul Sijben, and John J. Stanaway Jr. Service-centric networks. Bell Labs Technical Journal, 3(3):98–115, Autumn 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Mitra:1998:VDT

[MMR98] Debasis Mitra, John A. Morrison, and K. G. Ramakrishnan. VPN DESIGNER: A tool for design of multiservice virtual private networks. Bell Labs Technical Journal, 3(4):15–31, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Marion:1991:MPM

[MR91] Edwin D. Marion and Eric J.

Riddleberger. Modular project management. AT&T Technical Journal, 70(2):49–62, 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Melliar-Smith:1997:KSI

[MSHT97] C. Mark Melliar-Smith, Douglas E. Haggan, and William W. Troutman. Key steps to the integrated circuit. Bell Labs Technical Journal, 2(4):15–28, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Moreland:1996:STD

[MSSU96] James P. Moreland, Eileen M.

Sieli, Mary Socratous, and Susan C. Udovic. Small-team development in a competitive environment. AT&T Technical Journal, 75(1):15–23, 1996. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Mitchell:1990:BPS

[MVC90] James W. Mitchell, Jorge Louis Valdes, and Gary Cadet. Benign precursors for semiconductor processing. AT&T Technical Journal, 69(6):101–112, November/December 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Mitra:1997:HVP

[MZ97] Debasis Mitra and Ilze Ziedins. Hierarchical virtual partitioning — algorithms for virtual private networking. Bell Labs Technical Journal, 2(2):68–81, Summer 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Narayanamurti:1991:UST

[NA91] Venkatesh Narayanamurti and Dan E. Arvizu. Using Sandia technologies to improve national competitiveness. AT&T Technical Journal, 70(6):2–9, November/December 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Nagel:1998:ESL

[Nag98] Menno H. Nagel. Environmental supply-line engineering: Eco-supplier development coupled to eco-design — a new approach. Bell Labs Technical Journal, 3(2):109–123, Summer 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Netravali:1997:ISS

[Net97] Arun N. Netravali. The impact of solid-state electronics on computing and communications. Bell Labs Technical Journal, 2(4):126–154, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Narayanaswamy:1998:UDD

[NHK98] Shankar Narayanaswamy, Jianying Hu, and Ramanujan S. Kashi. Using data on digital cellular and PCS voice networks. Bell Labs Technical Journal, 3 (2):58–75, Summer 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Nissenbaum:1999:DHN

[Nis99]  Henry D. Nissenbaum.  A

directory-hosted network management information model: A CIM/DEN implementation. Bell Labs Technical Journal, 4 (4):55–74, October/December 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Nordin:1993:SDP

[NL93] Ronald A. Nordin and Anthony F. J. Levi. A system design perspective on optical interconnection technology. AT&T Technical Journal, 72(5):37–49, September/October 1993. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Noh:1998:ASS

[Noh98] Tai H. Noh. ATM scenarios for SDH/SONET networks. Bell Labs Technical Journal, 3 (1):81–93, Spring 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Odlyzko:1994:PKC

[Odl94] Andrew M. Odlyzko. Public key cryptography. AT&T Technical Journal, 73(5):17–23, September/October 1994. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (elec-

                                        tronic). URL http://www.
                                        research.att.com/~amo/doc/
                                        arch/public.key.crypto.pdf;
                                         http://www.research.att.
                                        com/~amo/doc/arch/public.
                                        key.crypto.ps;   http://www.
                                        research.att.com/~amo/doc/
                                        arch/public.key.crypto.tex.

Okrasinski:1995:EME

[OOM95] Thomas A. Okrasinski, Craig C. Onori, and Joseph M. Morabito. Energy management and the environment. AT&T Technical Journal, 74(6):44–53, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Opaluch:1993:TWI

[OT93] R. E. Opaluch and Yao-Chung Tsao. Ten ways to improve usability engineering-designing user interfaces for ease of use. AT&T Technical Journal, 72(3): 75–88, May/June 1993. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Patel:1990:MPC

[Pat90] C. Kumar N. Patel. Materials and processing: Core competencies and strategic resources. AT&T Technical Journal, 69(6): 2–8, November/December 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Parker:1999:UMV

[PBBV99] Martin F. Parker, Gordon R.

Brunson, W. J. Barnett Jr., and Gregory M. Vaudreuil. Unified messaging — a value-creating engine for next-generation network services. Bell Labs Technical Journal, 4(2):71–87, Summer 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Peercy:1991:SLS

[Pee91] Paul S. Peercy. Strained-layer semiconductor research and development at Sandia. AT&T Technical Journal, 70(6):49–58, November/December 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Penzias:1997:NFY

[Pen97] Arno A. Penzias. The next fifty years: Some likely impacts of solid-state technology. Bell Labs Technical Journal, 2(4): 155–168, Autumn 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Philen:1992:OFS

[PGGA^{+}92] D. L. Philen, M. Q. Garvin, J. Garcia-Arce, W. M. M. Botermans, and E. Gomez Chili. Optical fiber system design for TelMex. AT&T Technical Journal, 71(5):36–44, September/October 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Parker:1992:QIU

[PH92] T. Paul Parker and Gordon L. Harrison. Quality improvement using environmental stress testing. AT&T Technical Journal, 71(4):10–23, July/August 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Pittampalli:1999:TGC

[Pit99] Eshwar Pittampalli. Thirdgeneration CDMA wireless standards and harmonization. Bell Labs Technical Journal, 4 (3):6–18, Autumn 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Perdue:1990:AVP

[PLS90] R. J. Perdue, J. M. Leikness, and J. B. Sharp. AT&T voice processing system architectures. AT&T Technical Journal, 69 (5):52–60, September/October 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Pike:1999:SRA

[PR99] Rob Pike and Dennis M. Ritchie. The Styx(R) architecture for distributed systems. Bell Labs Technical Journal, 4 (2):146–152, Summer 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Porta:1996:DCP

[PSS96] Thomas F. La Porta, Krishan K. Sabnani, and Anil S. Sawkar. Distributed call processing for wireless mobile networks. Bell Labs Technical Journal, 1(2):127–142, Summer 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rabiner:1995:TVV

[Rab95] Lawrence R. Rabiner. Toward Vision 2001: voice and audio processing considerations. AT&T Technical Journal, 74(2): 4–13, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Russo:1991:INP

[RAMS^{+}91] Ernest G. Russo, M. Tawfik Abdel-Moneim, Linda L. Sand, Laura J. Shaw, Patricia D. Taska, and Ronald J. Wojcik. Intelligent network platforms in the US. AT&T Technical Journal, 70(3–4):26–43, March/ April 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Refi:1999:OFO

[Ref99] James J. Refi. Optical fibers for optical networking. Bell Labs Technical Journal, 4(1):

246–261, Spring 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rexing:1991:SPM

[Rex91] Gerald L. Rexing. Software project management. moving beyond project plans. AT&T Technical Journal, 70(2):40–48, 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Riggs:1996:STF

[RFS96] Victoria G. Riggs, Robert C. Fairfield, and Juan Segura. The speech transcoding frame 2000 for GSM cellular systems. Bell Labs Technical Journal, 1(2): 7–20, Summer 1996. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ryva:1991:CNE

[RHNS91] George J. Ryva, Brian T. Howard, John A. Newell, and John H. Sun. Corporate networking. evolution and architecture. AT&T Technical Journal, 70(5):2–13, September/October 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rodriguez-Moral:1997:LIF

[RM97] Antonio Rodriguez-Moral. LI-BRA — an integrated framework for type of service-based adaptive routing in the Internet and intranets. Bell Labs Technical Journal, 2(2):42–67, Summer 1997. CODEN BLTJFD.

ISSN 1089-7089 (print), 1538-7305 (electronic).

Roy:1997:LPT

[RMN97] Pradip K. Roy, Sailesh M. Merchant, and Arun K. Nanda. Layering processes and their applications in silicon microelectronics. Bell Labs Technical Journal, 2(2):234–245, Summer 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Roche:1997:HSD

[RMR^{+}97] Humberto J. La Roche, Stephen L.

Myers, James P. Runyon, Gerald S. Soloway, and David W. Spears. High-speed data services using the Switched Digital Broadband Access System. Bell Labs Technical Journal, 2 (2):188–202, Summer 1997. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ross:1997:FSA

[Ros97] Ian M. Ross. The foundation of the silicon age. Bell Labs Technical Journal, 2(4):3–14, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Rubinstein:1991:CNA

[RRW91] Charles B. Rubinstein, George J.

Ryva, and Peter S. Warwick. Corporate networking applications. AT&T Technical Journal, 70(5):27–35, September/October 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Reichmanis:1990:CLM

[RT90] Elsa Reichmanis and Larry F. Thompson. Challenges in lithographic materials and processes. AT&T Technical Journal, 69 (6):32–45, November/December 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Runge:1992:ULS

[Run92] Peter K. Runge. Undersea lightwave systems. AT&T Technical Journal, 71(1):5–13, January/February 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Rising:1998:IQP

[RW98] Linda S. Rising and Jack E. Watson. Improving quality and productivity in training: a new model for the hightech learning environment. Bell Labs Technical Journal, 3(1): 134–143, Spring 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Romascanu:1999:SMN

[RZ99] Dan Romascanu and Itai E. Zilbershtein. Switch monitoring — The new generation of monitoring for local area networks. Bell Labs Technical Journal, 4(4):42–54, October/December 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Schesser:1995:DRC

[SAES95] Joel Schesser, Stuart M. Abbott, Robert L. Easton, and M. S. Stix. Design requirements for the current generation of undersea cable systems. AT&T Technical Journal, 74(1):16–32, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Silverio:1993:EGK

[SBH^{+}93] Vincent J. Silverio, Sandra J.

Brown, Warren B. Houck, Giorgio Marinzoli, and Stacey D. Sern. Evolution of global key-system platforms. AT&T Technical Journal, 72(2):6–16, March/April 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Shankar:1999:MMC

[SC99] Ravi V. Shankar and Ashish K.

Chatterjee. Managing the management communications network in optical transport systems. Bell Labs Technical Journal, 4(4):155–170, October/December 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Scheer:1995:IUS

[Sch95]  Randall J. Scheer.  Inter-

nationalizing UNIX software projects. AT&T Technical Journal, 74(3):85–94, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Schultz:1998:AAD

[Sch98] Thomas A. Schultz. ADEPT — the Advanced Database Environment for Planning and Tracking. Bell Labs Technical Journal, 3(3):3–9, Autumn 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Scribner:1994:OSE

[Scr94]  N. Scribner.  The Opera-

tions System Evolution Laboratory. AT&T Technical Journal, 73(4):29–34, 1994. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Siller:1990:AMB

[SD90] Curtis A. Siller, Jr. and Walter Debus, Jr. Application of memory-based digital filtering to high-speed digital radio. AT&T Technical Journal, 69(5): 100–115, September/October 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Seifert:1995:ATE

[Sei95] Laurence C. Seifert. AT&T technology and the environment. AT&T Technical Journal, 74(6):4–7, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sharp:1994:NSH

[SEKS94] Ronald L. Sharp, Steven R. Eisen, W. E. Kleppinger, and Mark E. Smith. Network security in a heterogeneous environment. AT&T Technical Journal, 73(5):52–60, September/October 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Stulen:1991:DSX

[SF91] Richard H. Stulen and Richard R.

Freeman. Developing a soft Xray projection lithography tool. AT&T Technical Journal, 70 (6):37–48, November/December 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Spencer:1999:DRE

[SH99] Douglas A. Spencer and Kenneth W. Howard. DEFIN-ITY(R) enterprise communications server ATM integration. Bell Labs Technical Journal, 4 (2):21–42, Summer 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sieli:1991:MPP

[Sie91] Eileen M. Sieli. Managing a project as a process. AT&T Technical Journal, 70(2):33–39, 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sipress:1995:UCT

[Sip95] Jack M. Sipress. Undersea communications technology. AT&T Technical Journal, 74(1):4–7, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sable:1991:IND

[SK91] Edward G. Sable and Herbert W. Kettler. Intelligent network directions. AT&T Technical Journal, 70(3–4):2– 10, March/April 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sobrinho:1996:RTT

[SK96] João L. Sobrinho and A. S. Krishnakumar. Real-time traffic over the IEEE 802.11 medium access control layer. Bell Labs Technical Journal, 1(2): 172–187, Summer 1996. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Schmidt:1991:IEC

[SKS91] Judith L. Schmidt, Thaddeus J. Kowalski, and David S. Smull. An incremental environment for computer-aided design tools. AT&T Technical Journal, 70 (1):101–110, January/February 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Stafford:1995:UNR

[SMS95] E. K. Stafford, J. Mariano, and M. M. Sanders. Undersea nonrepeatered technologies, challenges, and products. AT&T Technical Journal, 74(1):47–59, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sproat:1995:TSS

[SO95] R. W. Sproat and J. P. Olive. Text-to speech synthesis. AT&T Technical Journal, 74(2):35–44, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sproat:1998:MTS

[Spr98] Richard William Sproat, editor.

Multilingual text-to-speech synthesis: the Bell Labs approach. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998. ISBN 0-7923-8027-

4. xxvi + 300 pp. LCCN TK7882.S65 M86 1998. URL

http://www.loc.gov/catdir/
enhancements/fy0821/97031463-
d.html;   http://www.loc.
gov/catdir/enhancements/fy0821/
97031463-t.html. With a fore-

word by Louis Pols.

Schulzrinne:1998:SIP

[SR98] Henning G. Schulzrinne and Jonathan D. Rosenberg. The Session Initiation Protocol: Providing advanced telephony services across the Internet. Bell Labs Technical Journal, 3(4): 144–160, October/December 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Schmalzried:1991:CVT

[SS91] T. E. Schmalzried and D. S.

Stewart. Characterizing voice transmission performance for evolving business networks. AT&T Technical Journal, 70 (5):69–80, September/October 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sansone:1992:ITM

[SS92] Frank P. Sansone and Howard M. Singer. Improving time to market in consumer products. AT&T Technical Journal, 71(3): 66–72, May/June 1992. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sijben:1998:BGI

[SS98] Paul G. A. Sijben and Louise F. A. Spergel. Bridging the gap to IP telephony. Bell Labs Technical Journal, 3(4):192–207, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Sherman:1994:SNA

[SSM94] S. A. Sherman, R. Skibo, and R. S. Murray. Secure network access using multiple applications of AT&T’s Smart Card. AT&T Technical Journal, 73 (5):61–72, September/October 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Seshadri:1993:ATM

[SSW93] Nambi Seshadri, Carl-Erik W. Sundberg, and Vijitha Weerackody. Advanced techniques for modulation, error correction, channel equalization, and diversity. AT&T Technical Journal, 72(4):48–63, July/August 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sun:1999:OFA

[SSZS99] Yan Sun, Atul K. Srivastava, Jianhui Zhou, and James W. Sulhoff. Optical fiber amplifiers for WDM optical networks. Bell Labs Technical Journal, 4 (1):187–206, Spring 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Stanzione:1994:OSP

[Sta94] Dan Stanzione. Operations systems: a perspective. AT&T Technical Journal, 73(4):4–13, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Subrahmanyam:1991:ASM

[Sub91] Pasupathi A. Subrahmanyam.

Automated synthesis of mixedmode (asynchronous and synchronous) systems. AT&T Technical Journal, 70(1):111– 124, January/February 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sulgrove:1996:SSP

[Sul96] Robert N. Sulgrove. Scoping software projects. AT&T Technical Journal, 75(1):35–45, 1996. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Sykes:1992:GIN

[SV92] Jack S. Sykes and Eldred J.

Visser. Globalization of intelligent network services. AT&T Technical Journal, 71(5):6–12, September/October 1992. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Szurkowski:1995:IMS

[SW95] Edward S. Szurkowski and Jack C. Warner. Interactive multimedia services for consumers and businesses. AT&T Technical Journal, 74(5):117– 129, September/October 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Stevens:1999:PBM

[SW99] Mark L. Stevens and Walter J. Weiss. Policy-based management for IP networks. Bell Labs Technical Journal, 4(4):75–94, October/December 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Torabi:1998:TGM

[TB98] Mohammad Torabi and Rolfe E. Buhrke. Third-generation mobile telecommunications and virtual home environment: A prioritization analysis. Bell Labs Technical Journal, 3(3): 50–62, Autumn 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Trotter:1998:LSD

[TCG^{+}98] John A. Trotter, Mark R. Cravatts, Gary L. Griffith, Paul Krzyzanowski, Wilfred E. Lehder Jr., Steve J. Nossen, and Cormac J. Sreenan. Lowspeed data over the PWT wireless phone system. Bell Labs Technical Journal, 3(1):19–29, Spring 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Tebbs:1999:RTI

[Teb99] Robert G. Tebbs. Real-time IP facsimile: Protocol and gateway requirements. Bell Labs Technical Journal, 4(2):128–145, Summer 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Teger:1995:MVR

[Teg95] Sandra L. Teger. Multimedia — from vision to reality. AT&T Technical Journal, 74(5):4–13, September/October 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Toth:1991:MAG

[Tot91] Michael S. Toth. MAKCAL: An application generator for advice. AT&T Technical Journal, 70(1): 9–20, January/February 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Townes:1999:HLH

[Tow99]  Charles H. Townes.  How

the Laser Happened: Adventures of a Scientist. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-512268-2 (hardcover), 0-19-515376-6, 1-280-47085-2. 200 pp. LCCN QC687.2 .T68 1999. URL

                                        http://www.loc.gov/catdir/
                                        enhancements/fy0603/98022216-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/fy0603/
                                        98022216-t.html;   http://
                                        www.loc.gov/catdir/enhancements/
                                        fy0724/98022216-b.html.

Thompson:1995:SCG

[TSS95] Robin J. Thompson, J. J. Schiet, and Steven L. Stevens. 5ESS-2000 switch cellular gateway. AT&T Technical Journal, 74(4):85–93, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Tuch:1993:DWI

[Tuc93]  Bruce Tuch.  Development

of WaveLAN, an ISM band wireless LAN. AT&T Technical Journal, 72(4):27–37, July/August 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Tull:1992:QPI

[Tul92] Bernie D. Tull, Jr. Quality and productivity in injection mold tool construction: A view from manufacturing. AT&T Technical Journal, 71(4):61–68, July/August 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Townsend:1995:UTB

[TWN95] Richard L. Townsend, Jr., Jean-Jacques Werner, and Mai-Huong Nguyen. Using technology to bring ATM to the desktop. AT&T Technical Journal, 74(4):25–37, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Tarallo:1999:O

[TZ99] Joseph A. Tarallo and George I.

Zysman. Overview. Bell Labs Technical Journal, 4(3): 3–5, Autumn 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Ufongene:1999:MCA

[Ufo99] Charles Mmaduka Ufongene. A model for the cost analysis of wireless access architectures. Bell Labs Technical Journal, 4 (3):134–154, Autumn 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Uyar:1990:AVI

[ULS90] M. U. Uyar, A. Lapone, and K. K. Sabnani. Algorithmic verification of ISDN network layer protocol. AT&T Technical Journal, 69(1):17–31, 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

vandenBroecke:1997:UDP

[vdBC97] Just A. van den Broecke and James O. Coplien. Using design patterns to build a framework for multimedia networking. Bell Labs Technical Journal, 2(1):166–187, Spring 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Veeraraghavan:1997:CCA

[Vee97] Malathi Veeraraghavan. Connection control in ATM networks. Bell Labs Technical Journal, 2(1):48–64, Spring 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Vehse:1991:MRC

[VNB91] Robert C. Vehse, Stephen F. Nygren, and A. Duane Butherus. Managing an R&D contract with the government. AT&T Technical Journal, 70(2):84–98, 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Verma:1990:NAS

[VPR^{+}90] Mahendra K. Verma, Dimitrios P. Prezas, Thomas L. Russell, Jr., Mostafa Hasham Sherif, and Reed Thorkildsen. Novel applications of speech processing in AT&T network systems products. AT&T Technical Journal, 69(5):77–86, September/October 1990. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

vanSoest:1997:BTW

[vSV97] Tim van Soest and René Verkuil. The best of three worlds — Building a distributed application as a step toward business objects. Bell Labs Technical Journal, 2(1):200– 219, Spring 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Walker:1990:DOF

[Wal90] Kenneth L. Walker. Directions in optical fibers. AT&T Technical Journal, 69(6):92– 100, November/December 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Wasser:1992:OD

[Was92] Daniel J. Wasser. Optical datalinks. AT&T Technical Journal, 71(1):46–52, January/February 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Wyatt:1991:EGI

[WBE^{+}91] George Y. Wyatt, Alvin Barshefsky, Robert V. Epley, Marc P. Kaplan, and Krish P. Krishnan. The evolution of global intelligent network architecture. AT&T Technical Journal, 70 (3–4):11–25, March/April 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Wei:1999:ECO

[WCW99] Leping Wei, Yi Chen, and Gerald G. Wong. The evolution of China’s optical fiber networks. Bell Labs Technical Journal, 4 (1):125–144, Spring 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wenger:1992:ASA

[WDM^{+}92] G. M. Wenger, D. A. Dickinson, G. C. Munie, G. P. Tashjian, and R. G. Furrow. Aqueous and semi-aqueous cleaning processes. AT&T Technical Journal, 71(2):45–51, 1992. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Weiss:1998:QDS

[Wei98] Walter Weiss. QoS with differentiated services. Bell Labs Technical Journal, 3(4):48–62, October/December 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Whang:1998:VPT

[WFK^{+}98] Kevin K. Whang, William Fink, Bala Krishnamurthy, I-Far Lin, David M. Rouse, and Henrik V. Sorensen. Voice over PacketStar^{T M} gateway solution for service provider networks. Bell Labs Technical Journal, 3(4):103–123, October/December 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wattenbarger:1993:SCA

[WGHL93] Blake L. Wattenbarger, Roger B. Garberg, Edward S. Halpern, and Barry L. Lively. Serving customers with automatic speech recognition – humanfactors issues. AT&T Technical Journal, 72(3):28–41, May/ June 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Whetzel:1996:IWW

[Whe96] J. K. Whetzel. Integrating the World Wide Web and database technology. AT&T Technical Journal, 75(2):38–46, 1996. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Whang:1999:NEM

[WJC99] Kevin K. W. Whang, Rajnish Jain, and Paul Chan. Network element management architecture for the 7 R/E^{T M} Trunk Access Gateway. Bell Labs Technical Journal, 4(4):204–224, October/December 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Workman:1991:IAA

[WKS^{+}91] Alexandra M. Workman, Murthy V.

Kolipakam, Janis B. Sharpless, Vilma Stoss, and Hans van der Veer. International applications of AT&T’s intelligent network platforms. AT&T Technical Journal, 70(3–4):44–57, March/ April 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Wang:1998:WVI

[WMGL98] Jin Wang, Peter J. McCann, Patvardhana B. Gorrepati, and Chung-Zin Liu. Wireless voiceover-IP and implications for third-generation network design. Bell Labs Technical Journal, 3(3):79–97, Autumn 1998. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Wilpon:1990:SRL

[WMRG90] Jay G. Wilpon, Rajendra P. Mikkilineni, David B. Roe, and Sedat Gokcen. Speech recognition: From the laboratory to the real world. AT&T Technical Journal, 69(5):14–24, September/October 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Wightman:1995:DEA

[WMS95] Glenn C. Wightman, Gregory C. Munie, and Bryan K. Stolte. Design for environment attributes of the AT&T 5ESS switch. AT&T Technical Journal, 74(6):61–71, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Wilson:1994:CPD

[WN94] Randall J. Wilson and John J. Mac Namara. Cellular and PCS digital-mobile-services switching center. AT&T Technical Journal, 73(6):55–61, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Weerackody:1996:TJC

[WPE96] Vijitha Weerackody, Christine Podilchuk, and Anthony Estrella. Transmission of JPEGcoded images over wireless channels. Bell Labs Technical Journal, 1(2):111–126, Summer 1996. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Weissman:1991:ECM

[WS91] S. H. Weissman and J. C. Sekutowski. Environmentally conscious manufacturing: a technology for the nineties. AT&T Technical Journal, 70(6):23–30, November/December 1991. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Wu:1991:PTI

[Wu91] E. Wu. PEST: a tool for implementing pseudo-exhaustive self-test. AT&T Technical Journal, 70(1):87–100, January/February 1991. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Wilson:1999:FFF

[WWS^{+}99] Gordon C. Wilson, Thomas H.

Wood, J. Anthony Stiles, Robert D. Feldman, Jean-Marc P. Delavaux, Thomas H. Daugherty, and Peter D. Magill. FiberVista: An FTTH or FTTC system delivering broadband data and CATV services. Bell Labs Technical Journal, 4 (1):300–322, Spring 1999. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Warwick:1993:TGS

[WYKO93] Collin A. Warwick, Ran-Hong Yan, Young Ock Kim, and Abbas Ourmazd. Toward gigascale silicon integrated circuits. AT&T Technical Journal, 72 (5):50–58, September/October 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Xu:1997:RST

[Xu97] Kevin Houzhi Xu. Reliable stream transmission protocols in mobile computing environments. Bell Labs Technical Journal, 2(3):152–163, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Yau:1994:BST

[YBC^{+}94] C. W. Yau, J. Beausang, F. E. Crane III, N. T. Jarwala, and R. E. Tulloss. Boundary-scan testing for electronic subassemblies and systems. AT&T Technical Journal, 73(2):40–48, 1994. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Yu:1997:ESS

[YBH97] Weider D. Yu, Alvin Barshefsky, and Steel T. Huang. An empirical study of software faults preventable at a personal level in a very large software development environment. Bell Labs Technical Journal, 2(3):221–232, Autumn 1997. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Yoder:1994:AAS

[YFF94] Keith W. Yoder, Gijs Franken, and Alan W. Fulton. AT&T ATM switching-network products. AT&T Technical Journal, 73(6):48–54, 1994. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Yochum:1996:IHR

[YLB96] D. S. Yochum, E. P. Laws, and G. K. Barlow. An integrated human resources approach to moving information technology professionals toward best in class. AT&T Technical Journal, 75(1):46–53, 1996. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Yeung:1995:QPU

[YS95] Man Keung Yeung and Tanjore K. Srinivas. Quick prototyping using JAM* for the customized FT-2 project. AT&T Technical Journal, 74(2):102– 110, ???? 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Yu:1990:SPM

[YSH90] W. D. Yu, P. Smith, and S. T. Huang. Software productivity measurements. AT&T Technical Journal, 69(3):110– 120, 1990. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Yu:1998:SFP

[Yu98] Weider D. Yu. A software fault prevention approach in coding and root cause analysis. Bell Labs Technical Journal, 3 (2):3–21, Summer 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zyskind:1992:EDF

[ZGSD92] John L. Zyskind, C. Randy Giles, Jay R. Simpson, and David J. DiGiovanni. Erbiumdoped fiber amplifiers and the next generation of lightwave systems. AT&T Technical Journal, 71(1):53–62, January/February 1992. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Zimmerman:1993:TMM

[Zim93] Michael A. Zimmerman. The technology of molded multichip modules. AT&T Technical Journal, 72(5):73–83, September/October 1993. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Zsakany:1995:AUC

[ZMRR95] J. C. Zsakany, N. W. Marshall, J. M. Roberts, and D. G. Ross. The application of undersea cable systems in global networking. AT&T Technical Journal, 74(1):8–15, 1995. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Zeger:1999:CCA

[ZN99] Linda M. Zeger and Mark E. Newbury. CDMA capacity with added carriers in a cellular network. Bell Labs Technical Journal, 4(3):104–119, Autumn 1999. CODEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zysman:1996:TWW

[ZTL96] George I. Zysman, Reed Thorkildsen, and David Y. Lee. Twoway wireless broadband access. Bell Labs Technical Journal, 1 (1):115–129, Spring 1996. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).

Zhu:1998:CQV

[ZYB98] Wenwu Zhu, Kyeong Ho Yang, and Marc J. Beacken. CIFto-QCIF video bitstream downconversion in the DCT domain. Bell Labs Technical Journal, 3 (3):21–29, Autumn 1998. CO-DEN BLTJFD. ISSN 1089-7089 (print), 1538-7305 (electronic).